### v1

# On the Fragility of Active Learners

Abstract
Active learning (AL) techniques aim to maximally utilize a labeling budget by iteratively selecting instances that are most likely to improve prediction accuracy. However, their benefit compared to random sampling has not been consistent across various setups, e.g., different datasets, classifiers. In this empirical study, we examine how a combination of different factors might obscure any gains from an AL technique.
Focusing on text classification, we rigorously evaluate AL techniques over around $1000$ experiments that vary wrt the dataset, batch size, text representation and the classifier. We show that AL is only effective in a narrow set of circumstances. We also address the problem of using metrics that are better aligned with real world expectations.
The impact of this study is in its insights for a practitioner: (a) the choice of text representation and classifier is as important as that of an AL technique, (b) choice of the right metric is critical in assessment of the latter, and, finally, (c) reported AL results must be holistically interpreted, accounting for variables other than just the query strategy.
Keywords:
First keyword Second keyword Another keyword.
1 Introduction
Within a supervised learning setup, Active Learning (AL) techniques use a Query Strategy (QS) to identify an unlabeled set of instances which is optimal in the following sense: if labelled and added to the training data, they lead to the greatest improvement in model accuracy, relative to any other same-sized set. In cases where labelling is expensive, using AL is cost-efficient compared to random sampling, because the model reaches greater accuracy with a smaller number of labelled instances. AL techniques can be both domain-agnostic, e.g., uncertainty-based sampling [16], or domain specific, e.g., Contrastive Active Learning [21] for Natural Language Processing (NLP).
Our focus here is AL for text classification. Today, a practitioner has easy access to a variety of powerful classifiers via packages such as scikit-learn [25] and Hugging Face [34], and text representations, such as Universal Sentence Encoding (USE) [3] and MPNet [31]. Within that context, we ask these questions: is the choice of AL algorithm still as important? Do AL techniques exhibit strong performance over random sampling with any combination of text representation and a classifier?
In this empirical work, we answer this question in the negative: AL techniques can be fragile, i.e., they are heavily influenced by other variables. This study is valuable for the following reasons:
  1. 1.
    In practice, where data labeling is costly, should one just spend time seeking a good AL algorithm? Or should they cast their net of design choices wider to take into account the classifier and representation as well?
  2. 2.
    An important part of assessing a technique is how we measure, i.e., metrics used. We discuss this aspect too, e.g., rank-based metrics such as the Wilcoxon test or Mean Rank ignore how large is the accuracy gap wrt random sampling.
    The practical implication of this is that a small gap might not justify the cost of an AL workflow, given that random sampling is trivial to use in most programming languages.
  3. 3.
    Finally, this cautions us as consumers of research that an AL technique is not just the QS; reported results must be carefully understood to be situated within, and affected by, other significant experiment variables such as the classifier, representation, batch size, etc.
To our knowledge, this is the first such comprehensive study of AL algorithms for the task of text classification. The rest of the paper is organized as follows:
2 Previous Work
The field of AL has seen a steady stream of novel contributions over the years, e.g., [16, 12, 29, 7]. Being a hard problem - identifying important instances having only seen a few examples from the instance-label distribution - it also seems to suffer from real-world challenges, e.g., wrt variations in dataset, batch sizes, classifier used, which has been pointed out in the literature [1, 35, 18, 20].
In the domain of NLP, the problem is compounded by the fact that text representations may vary widely. This is pertinent since a lot of rich representations have been proposed in the past few years, e.g., word2vec [22], Universal Sentence Encodings (USE) [3], Backpacks [14], BERT [10], MiniLM [33], MPNet [31]. We are not aware of any work that studies AL’s effectiveness in context of these relatively modern representations. Hence, this work constitutes a useful addition to the line of work that draws attention to challenges of AL deployment.
3 Batch Active Learning - Overview
In this work, we study AL techniques for the task of text classification, in the batch AL setting. To ensure no ambiguity, we provide pseudo-code for this AL setting in Algorithm 1.
Input: Unlabeled data $X_{U}$, test data $(X_{test},Y_{test})$, query strategy $\mathcal{Q}$, seed set selection strategy $\mathcal{A}$, search space $\Theta$ for model $\mathcal{M}$, seed size $s$, batch size $b$, number of iterations $T$, metric $\mathcal{V}$
Result: Scores on test set at various iterations $\{(\mathcal{V}_{0},0),(\mathcal{V}_{1},1),...,(\mathcal{V}_{T},T)\}$
$result\leftarrow\{\}$ // to be returned
1 $X_{L,0},X_{U,0}\leftarrow\mathcal{A}(X_{U},s)$
2 $(X_{L,0},Y_{L,0})\leftarrow\text{obtain labels for }X_{L,0}$
$M_{0}\leftarrow\operatorname*{arg\,max}_{\theta\in\Theta}M_{\theta}((X_{L,0},Y_{L,0}))$ // both model selection and calibration are performed
3 $\mathcal{V}_{0}\leftarrow\mathcal{V}(M_{0}(X_{test}),Y_{test})$
4 $result\leftarrow result\cup\{(\mathcal{V}_{0},0)\}$
5 for $t\leftarrow 1$ to $T$ do
6       $X_{L,t}^{new},X_{U,t}\leftarrow\mathcal{Q}(M_{t-1},X_{U,t-1},(X_{L,t-1},Y_{L,t-1}),b)$
7       $(X_{L,t}^{new},Y_{L,t}^{new})\leftarrow\text{obtain labels for }X_{L,t}^{new}$
8       $(X_{L,t},Y_{L,t})\leftarrow\text{add }(X_{L,t}^{new},Y_{L,t}^{new})\text{ to previous labeled pool }(X_{L,t-1},y_{L,t-1})$
       $M_{t}\leftarrow\operatorname*{arg\,max}_{\theta\in\Theta}M_{\theta}((X_{L,t},Y_{L,t}))$ // both model selection and calibration are performed
9       $\mathcal{V}_{t}\leftarrow\mathcal{V}(M_{t}(X_{test}),Y_{test})$
10       $result\leftarrow result\cup\{(\mathcal{V}_{t},t)\}$
11
12 end for
return $result$
Algorithm 1 Batch Active Learning.
At a high-level, at every AL iteration $1\leq t\leq T$, we use a query strategy $\mathcal{Q}$ to select a $b$-sized batch of instances from the unlabeled pool of data (line 8). We obtain labels for this set (line 9) and add it to the existing pool of labeled data (line 10). We then train a model $M_{t}$ over this data (line 11). We emphasize that:
  1. 1.
    The model $M_{t}$ is obtained after performing model selection over its hyperparameter space $\Theta$, using grid-search against a validation set. The validation set is a label-stratified subset (a $20\%$ split) of the current labeled set; the rest is used for training.
  2. 2.
    The model is also calibrated [^1] A notable exception is in our use of the RoBERTa model, which already is well calibrated [9].. This is critical since query strategies $\mathcal{Q}$ often use the predicted class probabilities from $M_{t}$. We use Platt scaling [26, 24].
The process is initialized by selecting a seed set of size $s$ from the unlabeled data pool, using a strategy $\mathcal{A}$ (line 2). We use random selection for this step.
We also note that a “model” here might mean a combination of a text representation, e.g., word vectors, and a classifier, e.g., Random Forest; further detailed in Section 4.
It might seem unusual that we explicitly mention model selection and calibration, which should be part of the standard practice for building probabilistic classifiers. Unfortunately, in our experience multiple implementations/studies miss one or both of these steps.
4 Comparison Methodology
In our experiments, we vary classifiers, text representations (we often jointly refer to them as a prediction pipeline), batch size, seed size and, of course, query strategies. These combinations are listed below:
  1. 1.
    Prediction pipeline: There are two categories of pipelines we use:
    1. (a)
      Separate representation and classifier: The representations used are USE [3], MPNet [31] and word vectors (we use the models provided by the spaCy library [15]). While USE and MPNet directly convert sentences into vectors, in the case of word vectors, we compute the average of the vectors of all words in a text to obtain its representation. For classification, we useRandom Forests (RF) [2] and Support Vector Machines [6] with a linear kernel - we’ll term the latter as “LinearSVC”.
      We use off-the-shelf representations and they are not fine-tuned on our data. Only the classifiers are trained on our data.
    2. (b)
      End-to-end classifier: This does not require a separate representation model. We use RoBERTa [17] (a variant of BERT). This is fine-tuned on the labeled data at each AL iteration.
    Hyperparameter search spaces are detailed in Section 0.A.4.2 of the Appendix. As noted in Section 3, model selection and calibration are performed during training of a prediction pipeline. The only exception is RoBERTa, which has been shown to be well-calibrated out of the box [9].
    The first category gives us $2\times 3=6$ combinations. Counting RoBERTa, we have $\boldsymbol{7}$ prediction pipelines in our study.
  2. 2.
    Query Strategy: we list these below, with the year of publication mentioned, to show our focus on contemporary techniques:
    1. (a)
      Random: the batch is selected uniformly at random. This forms our baseline.
    2. (b)
      Margin [^2] Also referred to as Smallest Margin or Breaking Ties. [27] (2001): this selects instances with the smallest differences between the confidence of the most likely and the second-most likely predicted (by the current classifier [^3] Note that in reference to Algorithm 1, at iteration $t$, the query strategy $\mathcal{Q}$ uses model $M_{t-1}$.) classes. Despite being a relatively old technique, it continues to be competitive [28].
    3. (c)
      Contrastive Active Learning (CAL) [21] (2021): chooses instances whose predicted class-probability distribution is the most different (based on KL divergence) from those of their $k$-nearest neighbors. This is similar to another work [23], where such conflicts are detected using the explanation space produced by XAI techniques.
    4. (d)
      Discriminative Active Learning (DAL) [13, 11] (2019): a binary classifier (a feedforward neural network) is constructed to discriminate between labeled and unlabeled data, and then selects unlabeled instances with the greatest predicted probability of being unlabeled. This picks examples that are most different from the labeled instances in this classifier’s representation space. While the original work [13] only considers image datasets, a separate study shows its efficacy on text [11].
    5. (e)
      Representative Errors for Active Learning (REAL) [5] (2023): identifies clusters in the unlabeled pool and assigns the majority predicted label as a “pseudo-label” to all points in it. Instances are then sampled whose predictions differ from the pseudo-label. The extent of disagreement and cluster size are factored into the sampling step
    We use a total of $\boldsymbol{5}$ query strategies.
  3. 3.
    Datasets: we use $\boldsymbol{5}$ standard datasets: agnews, sst-2, imdb, pubmed and dbpedia-5 (a $5$-label version of the standard dbpedia dataset that we created). These are detailed in Table 2 in the Appendix.
  4. 4.
    Batch and Seed sizes: We use batch and seed size combinations of $(200,200)$ and $(500,500)$. This is a total of $\boldsymbol{2}$ combinations.
  5. 5.
    Trials: For statistical significance, we run $\boldsymbol{3}$ trials for each combination of the above settings.
Other settings: The accuracy metric we use is the F1 (macro) score, since it prevents performance on a dominant class from overwhelming results. The size of the unlabeled pool is $20000$ at the start of each experiment. If the original dataset has more than than $20000$ instances, we extract a label-stratified sample, to retain the original class distribution. The size of the test set is $5000$ also a label-stratified sample from the corresponding test set of the original dataset.
We run an experiment till the size of the labeled set has grown to $5000$ instances. This which implies $T=(5000-200)/200=24$ iterations for the batch/seed size setting of $(200,200)$, and similarly $9$ iterations for the $(500,500)$ setting.
Total # experiments [^4] Our code is available here: https://anonymous.4open.science/r/On_the_Fragility_of_Active_Learners/README.md.: Accounting for the above settings, we have $7\times 5\times 5\times 2=350$ unique configurations. Accounting for the trials, we have a total of $350\times 3=\boldsymbol{1050}$ experiments. Reiterating something we had mentioned earlier: we follow the due process of performing model selection and calibration (excepting RoBERTa) in each AL iteration of each of these experiments.
5 Reproducibility of Benchmarked Techniques
Before we analyze our results, we would like to stress that our intention is not to suggest that the techniques we evaluate, e.g., REAL, CAL, DAL, don’t work. In the specific settings discussed in their respective papers, they most likely perform as reported. In the interest of fairness, we have conducted limited independent tests that confirm this - see Section 0.A.3 in the Appendix.
6 Results
We present our observations in this section. We begin with visualizing a cross-section of our experiments: Figure 1 shows results for the dataset agnews and batch/seed size of $(200,200)$.
Figure 1: F1 macro scores on the test set at each iteration, for the dataset agnews and batch size of $200$. The $x$-axes show size of the labeled data, the $y$-axes show the F1-macro scores on the test data. The corresponding plot for batch size = $500$ is shown as Figure 8 in the Appendix.
We introduce some notation to aid in the interpretation of our results. We think of the F1-macro score to be provided by this function:
$f(g,\mathcal{Q},\mathcal{D},b,s,N)$ (1)
Here,
  • $g$ denotes a prediction pipeline.
  • $\mathcal{Q}\in\{cal,dal,real,margin,random\}$ is a query strategy.
  • $\mathcal{D}\in\{imdb,agnews,dbpedia-5,pubmed,sst-2\}$ is a query strategy.
  • $b\in\{200,500\}$ is the batch size.
  • $s\in\{200,500\}$ is the seed size.
  • $N$ is the size of the labeled data. Here $N\leq 5000$.
[…]
We now try to answer some interesting questions based on our results. In most cases, we will average over various factors, especially datasets, since most AL techniques assume their query strategy may be used without any prior knowledge.
6.1 In what data regimes do query strategies most differ?
We immediately note in Figure 1 that learning curves vary quite a bit based on the prediction pipeline.
What broad trend(s) might we expect? F1-macro scores from different query strategies (for a given pipeline and dataset) should converge as we see more data due to at least two reasons: (a) the concept space would be eventually covered, (b) at later iterations, there is less of the unlabeled pool to choose from. Indeed, Figure 2 confirms this. We first compute variances in F1-macro scores for different pipeline/dataset combinations [^5] This step comes first since the accuracies obtained by a LinearSVC would be very different from those by RoBERTa. In this case, the inner quantity is conditioned over the dataset first, because we wanted to validate a specific property. In subsequent analyses, we will average over datasets, so as to mimic real-life use of AL, where no prior knowledge is assumed., and then plot their averages. In other words, a point $(x,y)$ for the batch/seed size $(200,200)$ curve denotes:
$y=\mathbb{E}_{g,\mathcal{D}}[Var[f|g=g^{\prime},\mathcal{D}=d^{\prime},N=x,b=200,s=200]]$ (2)
where $g^{\prime}$ and $d^{\prime}$ are a specific pipeline and dataset respectively.
Figure 2: Expectation over variance of F1-macro given a pipeline and dataset, plotted size of labeled data. See Section 6.1 for details. Note that the batch size doesn’t strongly affect trends.
6.2 Measuring effectiveness
A key reason to use AL is its effectiveness over random sampling. A common way to assess this is to use the Wilcoxon signed-rank test, e.g., [11], by comparing accuracy scores from these strategies, at various labeled data sizes. Unfortunately, this entirely ignores the magnitude of the difference in scores. This is a limitation of rank based metrics in general [^6] Another example is Mean Rank over AUC values of the learning curves.. See Figure 3, which shows learning curves for the random strategy, and query strategies $QS1$ and $QS2$. The latter two strategies have identical p-values of $9.77\times 10^{-4}$ for the Wilcoxon test (compared to random).
Figure 3: Hypothetical query strategies $QS1$ and $QS2$ shown against random. Note that p-value from the Wilcoxon test is identical for both. Discussed in Section 6.2.
In practice, one would clearly prefer $QS2$ over $QS1$, since it reaches higher accuracies earlier. In fact, considering the possibly non-zero cost of setting up an AL-based tagging system, a query strategy needs to be significantly better for it to be useful.
We recommend assessing the magnitude (relative to the random strategy) using Root Mean Square Error (RMSE) or the Area Under the (Learning) Curve (AUC).
RMSE vs AUC: the choice between them comes down to these factors: (a) RMSE doesn’t provide a direction of change (while AUC does), which may be addressed by pairing it with the Wilcoxon test, (b) AUC monotonically increases from left-to-right on the $x$-axis (unlike RMSE which is a mean), and so it must be carefully interpreted. Direct comparisons of AUC are most convenient over identical spans of the $x$-axis.
We apply these metrics to our results to obtain Figure 4, where we show both average p-values from the Wilcoxon test (on the left) and average RMSE scores (right subplot). Mathematically, for a given value on the $x$-axis, we compare the sequences:
$(\mathbb{E}_{g,\mathcal{D}}[f|\mathcal{Q}\in S_{qs},N=n,b=200,s=200])_{n\leq x}\text{ and}$
$(\mathbb{E}_{g,\mathcal{D}}[f|\mathcal{Q}=random,N=n,b=200,s=200])_{n\leq x}$ (3)
where $S_{qs}=\{cal,dal,real,margin\}$. Note that the $y$ values are cumulative, i.e., all sizes $n\leq x$ are considered. We prefer the RMSE here to normalize for the $x$-axis span.
For the Wilcoxon test, we have set the minimal sample size at $3$ points, which is why we see different starting points on the $x$-axes in Figure 4. This test is also performed to assess if the alternative, i.e., a non-random strategy, produces higher values than the random strategy.
Figure 4: The left plot shows averaged p-values from a Wilcoxon test, while the right plot shows averaged RMSE values. A consistently good query strategy should have a low p-value and high RMSE; none of the evaluate strategies fall in that category. Discussed in Section 6.2.
An ideal query strategy should have a low p-value and high RMSE. Unfortunately, no query strategy fulfills this criteria: while cal and dal have high RMSE values, they have poor p-values (the dal curve eclipses cal in the plot). And while real and margin have p-values that improve with labeled data, they exhibit small RMSE scores. Some things of note:
  1. 1.
    The fact the p-value scores of real and margin improve with size, is something we will encounter again (Figure 5).
  2. 2.
    The RMSE plot reinforces trends of the variance plot in Figure 2: we see higher RMSEs in the low data regime.
  3. 3.
    The batch/seed settings preserve the general trend.
6.3 Gain over random and margin strategies
We now ask if we actually do better than random sampling. We define the relative improvement $\delta$ between two scores as:
$\delta(treatment,control)=100\times\frac{(treatment-control)}{control}$ (4)
For a labeled dataset of size $(n_{1},n_{2}]$, batch/seed size $(b^{\prime},s^{\prime})$ and
$q^{\prime}\in\{cal,dal,real,margin\}$, we compute:
$\mathbb{E}_{g,\mathcal{D}}[\delta((f|\mathcal{Q}=q^{\prime},b=b^{\prime},s=s^{\prime},n_{1}<N\leq n_{2}]),\\ (f|\mathcal{Q}=random,b=b^{\prime},s=s^{\prime},n_{1}<N\leq n_{2}))]$ (5)
This is shown in Figure 5, where the labeled data size is broken up into multiple bins to distinguish patterns at different size regimes. We also separate the batch sizes (see legend). Given the negative scores, it seems that in general, random is either competitive or better!
At relatively higher data sizes, margin and real start to surpass random: which seems to suggest that (a) at low size regimes, where we expect the greatest impact from AL, their performance is actually poor, and (b) after gathering some data, these strategies are able to better exploit structure in the data (and margin seems surprisingly competitive in this regime). Also, recall that we had seen this behaviour of margin and real in the RMSE subplot of Figure 4.
Since behaviour across the two batch/seed size settings don’t seem to appreciably differ (we note this in Figures 2, 4 and 5), we will average over them in the next few plots.
Figure 5: Average relative improvement in F1-macro score over random, for different query strategies. Corresponds to Equation 5.
We can repeat the above exercise, to assess the effect of pipelines now, i.e., for a given data size bin $(n_{1},n_{2}]$ and prediction pipeline $g^{\prime}$, we now compute:
$\mathbb{E}_{\mathcal{Q}\in S_{qs},\mathcal{D},b,s}[\delta((f|g=g^{\prime},n_{1}<N\leq n_{2}]),\\ (f|g=g^{\prime},\mathcal{Q}=random,n_{1}<N\leq n_{2}))]$ (6)
where $S_{qs}=\{cal,dal,real,margin\}$. This is plotted in Figure 6. Here we note that, using RoBERTa gives us a better average performance over random at any size regime.
Figure 6: Average relative improvement in F1-macro score over random, for different pipelines. This corresponds to Equation 6.
We might want to know if for the promising case of RoBERTa relatively new techniques like dal, real and cal outperform margin. This is easy to plot with minor changes to the quantity in Equation 6:
$\mathbb{E}_{\mathcal{Q}\in S_{qs},\mathcal{D},b,s}[\delta((f|g=RoBERTa,n_{1}<N\leq n_{2}]),\\ (f|g=RoBERTa,\mathcal{Q}=margin,n_{1}<N\leq n_{2}))]$ (7)
where $S_{qs}=\{cal,dal,real\}$. The corresponding plot is shown in Figure 7. We see that aside from one data regime, $(1250-2500]$ margin remains quite competitive.
Figure 7: Average relative improvement in F1-macro score over margin for RoBERTa. This corresponds to Equation 7.
6.4 The best combinations
Finally, we ask what are the best combinations of pipelines and query strategies, at different size regimes? We measure this by the AUC per regime. Given a pipeline $g^{\prime}$, strategy $q^{\prime}$, and data sizes $n_{i}$ within a bin, the AUC [^7] Using AUC is reasonable since the $x$-axis is broken up into bins of roughly equal size (see discussion in Section 6.2), and we are not looking for comparison to the random strategy. is computed for the following sequence:
$(\mathbb{E}_{\mathcal{D},b,s}[f|g=g^{\prime},\mathcal{Q}=q^{\prime},N=n_{i}])_{n_{i}}$ (8)
Table 1 shows the top-$10$ combinations of pipelines and query strategies for different regimes. Not surprisingly, wrt pipelines, RoBERTa shows the strongest performance. Somewhat surprisingly, the next best pipeline is LinearSVC-MPNet (shortened to LinSVC_MPNet) and not a RF-based pipeline. Our guess as to why this is the case is that RF’s hyperparamṡearch space (detailed in Section 0.A.4.2) was not comprehensive enough, while MPNet transformed the data to be nearly linearly separable, which enabled the high-bias LinearSVC to quickly find a good optima. Sadly, no single query strategy dominates the QS column, and thus we are offered no insight as to what strategy pairs best with a pipeline. Somewhat noteworthy is the fact that the random strategy is quite competitive for the LinearSVC-MPNet pipeline, but not for RoBERTa.
Table 1: The top-$10$ combinations of pipelines and query strategies are shown (in two columns) for different labeled data size regimes. We show the AUC score for a combinations in size regime. See Section 6.4 for details.
train size pipeline QS AUC pipeline QS AUC
(0-1250] 1. RoBERTa dal 875.46 6. LinSVC_MPNet random 807.04
(0-1250] 2. RoBERTa real 873.64 7. LinSVC_MPNet real 794.69
(0-1250] 3. RoBERTa margin 873.36 8. LinSVC_MPNet margin 793.08
(0-1250] 4. RoBERTa cal 871.92 9. LinSVC_MPNet cal 790.42
(0-1250] 5. RoBERTa random 862.99 10. LinSVC_MPNet dal 789.60
(1250-2500] 1. RoBERTa cal 986.22 6. LinSVC_MPNet random 913.76
(1250-2500] 2. RoBERTa dal 984.40 7. LinSVC_MPNet real 907.98
(1250-2500] 3. RoBERTa real 981.07 8. LinSVC_MPNet margin 907.39
(1250-2500] 4. RoBERTa margin 976.07 9. LinSVC_MPNet dal 900.13
(1250-2500] 5. RoBERTa random 974.04 10. LinSVC_MPNet cal 899.12
(2500-3750] 1. RoBERTa margin 903.60 6. LinSVC_MPNet random 837.91
(2500-3750] 2. RoBERTa dal 900.93 7. LinSVC_MPNet margin 836.57
(2500-3750] 3. RoBERTa real 900.83 8. LinSVC_MPNet real 833.95
(2500-3750] 4. RoBERTa cal 897.16 9. LinSVC_MPNet dal 828.73
(2500-3750] 5. RoBERTa random 891.58 10. LinSVC_MPNet cal 827.24
(3750-5000] 1. RoBERTa real 1087.92 6. LinSVC_MPNet margin 1016.94
(3750-5000] 2. RoBERTa margin 1087.37 7. LinSVC_MPNet random 1010.88
(3750-5000] 3. RoBERTa dal 1085.06 8. LinSVC_MPNet real 1008.67
(3750-5000] 4. RoBERTa cal 1076.79 9. LinSVC_MPNet dal 1004.44
(3750-5000] 5. RoBERTa random 1075.20 10. LinSVC_MPNet cal 999.97
7 Summary and Conclusion
We first summarize our observations:
  1. 1.
    We noted in the dataset cross-section (Figure 1) of our experiments that the behavior of query strategies vary widely wrt pipelines. However, we observe the larger expected trend of reduction in the variance of accuracies produced by different strategies, with increase in the size of labeled data (Figure 2 in Section 2).
  2. 2.
    Wrt assessing the value of a query strategy, we noted that the Wilcoxon test only provides a partial picture by ignoring magnitudes (Figure 3 in Section 6.2). We might use the RMSE score to measure differences wrt the random strategy. Applied to our results, it shows that no query strategy exhibits the favorable properties of having both a low p-value and high RMSE score (Figure 4 in Section 6.2).
  3. 3.
    Our batch/seed size settings of $(200,200)$ and $(500,500)$ only seem to influence slight change in behaviors. We noted this in Figure 2, Figure 4 and Figure 5.
  4. 4.
    Most query strategies exhibit no significant gain over the random strategy (Figure 5), although at relatively higher data regimes we noted real and margin surpass random.
  5. 5.
    Among pipelines, RoBERTa exhibits the strongest performance (Figure 6) relative to random, but not wrt margin (Figure 7).
  6. 6.
    When we list the top-$10$ best performing combinations per data size regime (Table 1), we note that RoBERTa and LinearSVC-MPNet consistently do well, but unfortunately, their best performances are paired with different query strategies. The only (broad) takeaways are that while LinearSVC-MPNet works well with the random strategy, RoBERTa performs well with non-random strategies, but its not clear which specific one.
After extensive evaluation of different AL algorithms, we are forced to conclude that AL doesn’t reliably work across datasets, prediction pipelines and text representations. This study may be considered “limited” relative to real-world variances, e.g., many more choices of classifiers, datasets, which leads us to suspect that the true picture is probably more dismal. Of course, we see some associations in isolated pockets, such as RoBERTa consistently performs well, that are promising, but not enough (with what query strategy?) to derive a useful prescription. It’s known that query strategies require a “warm-start” labeled set, and this is probably why we observe the deferred performance of real and margin in Figure 5, but again, this does not translate to practical advice unless we can precisely identify a settings-specific warm-start size. It’d seem the only way to know if an AL technique would work in a setting is to find results reported for similar settings, and even this advice leaves open the question of inferring similarity between datasets. And if one can’t find such results, the only other way to know is by trying out multiple techniques - which of course is a chicken-and-egg problem when data-labeling is expensive!
How do we encourage the field of AL to be more utility-driven? We believe there are at least two ways: (a) embrace a broader discourse where the success of a technique needs to be tied to fundamental properties of datasets (e.g., topological features [4]) and predictors (e.g., VC dimension [32]), that are identifiable in an unsupervised manner in novel settings, and (b) on the empirical front, reported results need to account for an array of settings much more diverse than is customary today.
Appendix 0.A Appendix
0.A.1 Datasets
As an addendum to the methodology discussion in Section 4, we provide details of the datasets used in Table 2.
Table 2: Datasets used.
Dataset
# classes
Description
sst-2
2
Single sentences extracted from movie reviews with their sentiment label [30].
imdb
2
Movie reviews with corresponding sentiment label [19].
agnews
4
News articles with their topic category [36].
pubmed
4
Sentences in medical articles’ abstracts which are labeled with their role on the abstract [8].
dbpedia-5
5
A subset of dbpedia [36] which contains Wikipedia articles accompanied by a topic label. The original dataset’s instances are evenly distributed across $14$ classes. To form dbpedia-5, we use only the first $5$ classes: Company, EducationalInstitution, Artist, Athlete, OfficeHolder. This was done to reduce the training time of one-vs-all classifiers, e.g., LinearSVC.
0.A.2 Learning curves for agnews
This is the plot for the batch/seed size setting of $(500,500)$, for the dataset agnews. This is a companion to the plot shown in Figure 1 of Section 6.
Figure 8: F1 macro scores on the test set at each iteration, for the dataset agnews and batch size of $500$. The $x$-axes show size of the labeled data, the $y$-axes show the F1-macro scores on the test data.
0.A.3 Reproducibility Experiments
In all cases, we have attempted to replicate the original settings, e.g. same train/development/test data split, model type, seed/batch sizes, number of AL iterations as shown in Table 3. For CAL, REAL, we report the F1-macro scores on agnews, in which classes are evenly distributed, instead of the accuracy provided in the original papers. For DAL, we use the dataset cola [^8] https://nyu-mll.github.io/CoLA/ and utilise the Hugging Face library to finetune BERT (while the original work employs TensorFlow [^9] https://www.tensorflow.org/, but we use equivalent settings). Figure 9 shows a comparison between our results and the reported ones in these papers. Despite some minor differences in the setups, we observe that these AL methods work as described in their respective papers in these settings.
One significant difference between these settings compared to our methodology is the use of a predetermined labeled development set for all BERT/RoBERTa model finetuning. This set is relatively larger than the AL batch or seed size and is not part the labeled data available at each AL iteration. This is impractical in scenarios where AL is typically used: labeling is expensive. Moreover, in some cases, there is no model selection performed, which we remedy in our experiments (Section 3).
Figure 9: Comparison between published results and ours with the same settings for CAL, REAL, DAL.
Table 3: Settings for reproducibility experiments.
AL
Dataset
AL loop
Classifier & text representation
QS parameters
Metric
CAL
agnews
b=2280
s=1140
T = 7
BERT (bert-base-cased)
[CLS] at the last hidden layer
learning rate = 2e-5
train batch size = 16
# epochs = 3
sequence length = 128
warmup ratio = 0.1
# evaluations per epoch = 5
# neighbors=10
F1-macro
DAL
cola
b=50
s=100
T = 5
BERT (bert-base-uncased)
[CLS] at the pooled layer
learning rate = 5e-5
train batch size = 50
# epochs = 5
sequence length = 50
warmup ratio = 0
# evaluations per epoch = 1
Accuracy
REAL
agnews
b=150
s=100
T = 8
RoBERTa (roberta-base)
[CLS] at the last hidden layer
learning rate = 2e-5
train batch size = 8
# epochs = 4
sequence length = 96
warmup ratio = 0.1
# evaluations per epoch = 4
# clusters=25
F1-macro
0.A.4 Hyperparameters
0.A.4.1 Query Strategy (QS) hyperparameters
For each QS’s hypeparameters, we use the values recommended by the authors in corresponding papers. This means setting number of nearest neighbors in CAL to 10, number of clusters in DAL to 25, and keeping the same discriminative model in REAL.
0.A.4.2 Hyperparameters search for prediction pipelines
Table 4 shows the search space for hyperparameters we use for each classifier.
Table 4: Hyperparameters for each classifier in the prediction pipelines.
Classifier
Hyperparameters
RoBERTa
roberta-base
[CLS] at the last hidden layer
learning rate = {3e-5, 5e-5}
train batch size = 16
# epochs = {5, 10}
sequence length = 128
warmup ratio = 0.1
# evaluations per epoch = 5
LinearSVC
C = {0.001, 0.01, 0.1, 1, 10, 100, 1000}
class weight = balanced
RF
min samples leaf = {1, 5, 9}
# estimators = {5, 10, 20, 30, 40, 50}
max depth = {5, 10, 15, 20, 25, 30}
class weight = balanced
max features = sqrt

### v2

@@
 We run an experiment till the size of the labeled set has grown to $5000$ instances.
-This which implies $T=(5000-200)/200=24$ iterations for the batch/seed size setting of $(200,200)$, and similarly $9$ iterations for the $(500,500)$ setting.
+This implies $T=(5000-200)/200=24$ iterations for the batch/seed size setting of $(200,200)$, and similarly $9$ iterations for the $(500,500)$ setting.
 Total # experiments [^4] Our code is available here: https://anonymous.4open.science/r/On_the_Fragility_of_Active_Learners/README.md.: Accounting for the above settings, we have $7\times 5\times 5\times 2=350$ unique configurations.
@@
   • $\mathcal{Q}\in\{cal,dal,real,margin,random\}$ is a query strategy.
-  • $\mathcal{D}\in\{imdb,agnews,dbpedia-5,pubmed,sst-2\}$ is a query strategy.
+  • $\mathcal{D}\in\{imdb,agnews,dbpedia-5,pubmed,sst-2\}$ is a dataset.
   • $b\in\{200,500\}$ is the batch size.
@@
 Somewhat surprisingly, the next best pipeline is LinearSVC-MPNet (shortened to LinSVC_MPNet) and not a RF-based pipeline.
-Our guess as to why this is the case is that RF’s hyperparamṡearch space (detailed in Section 0.A.4.2) was not comprehensive enough, while MPNet transformed the data to be nearly linearly separable, which enabled the high-bias LinearSVC to quickly find a good optima.
+Our guess as to why this is the case is that RF’s hyperparam.
+search space (detailed in Section 0.A.4.2) was not comprehensive enough, while MPNet transformed the data to be nearly linearly separable, which enabled the high-bias LinearSVC to quickly find a good optima.
 Sadly, no single query strategy dominates the QS column, and thus we are offered no insight as to what strategy pairs best with a pipeline.
@@
 (0-1250] 1. RoBERTa dal 875.46 6. LinSVC_MPNet random 807.04
-(0-1250] 2. RoBERTa real 873.64 7. LinSVC_MPNet real 794.69
-(0-1250] 3. RoBERTa margin 873.36 8. LinSVC_MPNet margin 793.08
-(0-1250] 4. RoBERTa cal 871.92 9. LinSVC_MPNet cal 790.42
-(0-1250] 5. RoBERTa random 862.99 10. LinSVC_MPNet dal 789.60
+2. RoBERTa real 873.64 7. LinSVC_MPNet real 794.69
+3. RoBERTa margin 873.36 8. LinSVC_MPNet margin 793.08
+4. RoBERTa cal 871.92 9. LinSVC_MPNet cal 790.42
+5. RoBERTa random 862.99 10. LinSVC_MPNet dal 789.60
 (1250-2500] 1. RoBERTa cal 986.22 6. LinSVC_MPNet random 913.76
-(1250-2500] 2. RoBERTa dal 984.40 7. LinSVC_MPNet real 907.98
-(1250-2500] 3. RoBERTa real 981.07 8. LinSVC_MPNet margin 907.39
-(1250-2500] 4. RoBERTa margin 976.07 9. LinSVC_MPNet dal 900.13
-(1250-2500] 5. RoBERTa random 974.04 10. LinSVC_MPNet cal 899.12
+2. RoBERTa dal 984.40 7. LinSVC_MPNet real 907.98
+3. RoBERTa real 981.07 8. LinSVC_MPNet margin 907.39
+4. RoBERTa margin 976.07 9. LinSVC_MPNet dal 900.13
+5. RoBERTa random 974.04 10. LinSVC_MPNet cal 899.12
 (2500-3750] 1. RoBERTa margin 903.60 6. LinSVC_MPNet random 837.91
-(2500-3750] 2. RoBERTa dal 900.93 7. LinSVC_MPNet margin 836.57
-(2500-3750] 3. RoBERTa real 900.83 8. LinSVC_MPNet real 833.95
-(2500-3750] 4. RoBERTa cal 897.16 9. LinSVC_MPNet dal 828.73
-(2500-3750] 5. RoBERTa random 891.58 10. LinSVC_MPNet cal 827.24
+2. RoBERTa dal 900.93 7. LinSVC_MPNet margin 836.57
+3. RoBERTa real 900.83 8. LinSVC_MPNet real 833.95
+4. RoBERTa cal 897.16 9. LinSVC_MPNet dal 828.73
+5. RoBERTa random 891.58 10. LinSVC_MPNet cal 827.24
 (3750-5000] 1. RoBERTa real 1087.92 6. LinSVC_MPNet margin 1016.94
-(3750-5000] 2. RoBERTa margin 1087.37 7. LinSVC_MPNet random 1010.88
-(3750-5000] 3. RoBERTa dal 1085.06 8. LinSVC_MPNet real 1008.67
-(3750-5000] 4. RoBERTa cal 1076.79 9. LinSVC_MPNet dal 1004.44
-(3750-5000] 5. RoBERTa random 1075.20 10. LinSVC_MPNet cal 999.97
+2. RoBERTa margin 1087.37 7. LinSVC_MPNet random 1010.88
+3. RoBERTa dal 1085.06 8. LinSVC_MPNet real 1008.67
+4. RoBERTa cal 1076.79 9. LinSVC_MPNet dal 1004.44
+5. RoBERTa random 1075.20 10. LinSVC_MPNet cal 999.97
 7 Summary and Conclusion
@@
 For DAL, we use the dataset cola [^8] https://nyu-mll.github.io/CoLA/ and utilise the Hugging Face library to finetune BERT (while the original work employs TensorFlow [^9] https://www.tensorflow.org/, but we use equivalent settings).
-Figure 9 shows a comparison between our results and the reported ones in these papers.
+Figure 9 shows a comparison between our results and the reported ones in these papers [21, 5, 11] for CAL, REAL, DAL, respectively.
 Despite some minor differences in the setups, we observe that these AL methods work as described in their respective papers in these settings.
@@
 Moreover, in some cases, there is no model selection performed, which we remedy in our experiments (Section 3).
-Figure 9: Comparison between published results and ours with the same settings for CAL, REAL, DAL.
+Figure 9: Comparison between published results in [21, 5, 11] and ours with the same settings for CAL, REAL, DAL.
 Table 3: Settings for reproducibility experiments.

### v3

@@
 Table 2: Datasets used.
-Dataset
-# classes
-Description
-sst-2
-2
-Single sentences extracted from movie reviews with their sentiment label [30].
-imdb
-2
-Movie reviews with corresponding sentiment label [19].
-agnews
-4
-News articles with their topic category [36].
-pubmed
-4
-Sentences in medical articles’ abstracts which are labeled with their role on the abstract [8].
-dbpedia-5
-5
-A subset of dbpedia [36] which contains Wikipedia articles accompanied by a topic label.
+Dataset # classes Description
+sst-2 2 Single sentences extracted from movie reviews with their sentiment label [30].
+imdb 2 Movie reviews with corresponding sentiment label [19].
+agnews 4 News articles with their topic category [36].
+pubmed 4 Sentences in medical articles’ abstracts which are labeled with their role on the abstract [8].
+dbpedia-5 5 A subset of dbpedia [36] which contains Wikipedia articles accompanied by a topic label.
 The original dataset’s instances are evenly distributed across $14$ classes.
@@
 Table 3: Settings for reproducibility experiments.
-AL
-Dataset
-AL loop
-Classifier & text representation
-QS parameters
-Metric
-CAL
-agnews
-b=2280
+AL Dataset AL loop Classifier & text representation QS parameters Metric
+CAL agnews b=2280
 s=1140
@@
 # evaluations per epoch = 5
-# neighbors=10
-F1-macro
-DAL
-cola
-b=50
+# neighbors=10 F1-macro
+DAL cola b=50
 s=100
@@
 - Accuracy
-REAL
-agnews
-b=150
+REAL agnews b=150
 s=100
@@
 # evaluations per epoch = 4
-# clusters=25
-F1-macro
+# clusters=25 F1-macro
 0.A.4 Hyperparameters
@@
 Table 4: Hyperparameters for each classifier in the prediction pipelines.
-Classifier
-Hyperparameters
-RoBERTa
-roberta-base
+Classifier Hyperparameters
+RoBERTa roberta-base
 [CLS] at the last hidden layer
@@
 # evaluations per epoch = 5
-LinearSVC
-C = {0.001, 0.01, 0.1, 1, 10, 100, 1000}
+LinearSVC C = {0.001, 0.01, 0.1, 1, 10, 100, 1000}
 class weight = balanced
-RF
-min samples leaf = {1, 5, 9}
+RF min samples leaf = {1, 5, 9}
 # estimators = {5, 10, 20, 30, 40, 50}

### v5

@@
-On the Fragility of Active Learners
+On the Fragility of Active Learners for Text Classification
 Abstract
-Active learning (AL) techniques aim to maximally utilize a labeling budget by iteratively selecting instances that are most likely to improve prediction accuracy.
-However, their benefit compared to random sampling has not been consistent across various setups, e.g., different datasets, classifiers.
-In this empirical study, we examine how a combination of different factors might obscure any gains from an AL technique.
-Focusing on text classification, we rigorously evaluate AL techniques over around $1000$ experiments that vary wrt the dataset, batch size, text representation and the classifier.
-We show that AL is only effective in a narrow set of circumstances.
-We also address the problem of using metrics that are better aligned with real world expectations.
-The impact of this study is in its insights for a practitioner: (a) the choice of text representation and classifier is as important as that of an AL technique, (b) choice of the right metric is critical in assessment of the latter, and, finally, (c) reported AL results must be holistically interpreted, accounting for variables other than just the query strategy.
-Keywords:
-First keyword Second keyword Another keyword.
+Active learning (AL) techniques optimally utilize a labeling budget by iteratively selecting instances that are most valuable for learning.
+However, they lack “prerequisite checks”, i.e., there are no prescribed criteria to pick an AL algorithm best suited for a dataset.
+A practitioner must pick a technique they trust would beat random sampling, based on prior reported results, and hope that it is resilient to the many variables in their environment: dataset, labeling budget and prediction pipelines.
+The important questions then are: how often on average, do we expect any AL technique to reliably beat the computationally cheap and easy-to-implement strategy of random sampling?
+Does it at least make sense to use AL in an “Always ON” mode in a prediction pipeline, so that while it might not always help, it never under-performs random sampling?
+How much of a role does the prediction pipeline play in AL’s success?
+We examine these questions in detail for the task of text classification using pre-trained representations, which are ubiquitous today.
+Our primary contribution here is a rigorous evaluation of AL techniques, old and new, across setups that vary wrt datasets, text representations and classifiers.
+This unlocks multiple insights around warm-up times, i.e., number of labels before gains from AL are seen, viability of an “Always ON” mode and the relative significance of different factors.
+Additionally, we release a framework for rigorous benchmarking of AL techniques for text classification.
+On the Fragility of Active Learners for Text Classification
+Abhishek Ghose [24]7.ai abhishek.ghose.82@gmail.com                        Emma Thuong Nguyen [24]7.ai emma.ttn8@gmail.com
 1 Introduction
 Within a supervised learning setup, Active Learning (AL) techniques use a Query Strategy (QS) to identify an unlabeled set of instances which is optimal in the following sense: if labelled and added to the training data, they lead to the greatest improvement in model accuracy, relative to any other same-sized set.
-In cases where labelling is expensive, using AL is cost-efficient compared to random sampling, because the model reaches greater accuracy with a smaller number of labelled instances.
-AL techniques can be both domain-agnostic, e.g., uncertainty-based sampling [16], or domain specific, e.g., Contrastive Active Learning [21] for Natural Language Processing (NLP).
-Our focus here is AL for text classification.
-Today, a practitioner has easy access to a variety of powerful classifiers via packages such as scikit-learn [25] and Hugging Face [34], and text representations, such as Universal Sentence Encoding (USE) [3] and MPNet [31].
-Within that context, we ask these questions: is the choice of AL algorithm still as important?
-Do AL techniques exhibit strong performance over random sampling with any combination of text representation and a classifier?
-In this empirical work, we answer this question in the negative: AL techniques can be fragile, i.e., they are heavily influenced by other variables. This study is valuable for the following reasons:
+In cases where labelling is expensive, the value proposition of AL is that it is cost-efficient compared to random sampling, and a model reaches greater accuracy with a smaller number of labelled instances.
+In practice, an AL technique is selected based on the strength of prior reported results, i.e., there are no “prerequisite checks”: tests that one might perform on an unlabeled dataset, that help to select a technique suited to a problem [^1] We refer to this as the practitioner’s decision model and formalize it in §4.4..
+This trust extends to related decisions such as batch and seed sizes, as well as the hyperparameters (if any) of the AL technique since there is no way to empirically pick them: to compare with random sampling, or among techniques, labels are required.
+But if one had labels, they wouldn’t need AL!
+In this sense, the AL setup is unforgiving as one needs to make the optimal choice in one shot.
+This leads us to question the validity of the implicit but consequential assumption of task transfer.
+A related question is if it makes sense to use AL in an “Always ON” mode in a data labeling workflow; this is akin to asking if AL might perform worse than random sampling.
+We need to quantify both the frequency and magnitude of gains from AL, to be able to evaluate the cost of such pipeline; even simple AL techniques require a model to be evaluated over the unlabeled data pool, which can be expensive depending on the model complexity, size of the data pool and the latency allowed per AL iteration.
+To be clear, we don’t question if AL results are reproducible within the original setups they were reported in [^2] In the interest of fairness, we conducted limited reproduciblity tests for the AL techniques we benchmark here, and were able to replicate reported results - see §D.
+; but whether any of those gains carry forward to new setups, which is how AL is used in practice.
+We pick the area of text classification to investigate these concerns.
+The larger area of NLP has seen a rapid infusion of novel ideas of late.
+Today, a practitioner has easy access to a variety of powerful classifiers via packages such as scikit-learn Pedregosa et al.
+(2011), spaCy Honnibal et al.
+(2020) and Hugging Face Wolf et al.
+(2020), and text representations, such as Universal Sentence Encoding (USE) Cer et al.
+(2018), MiniLM Wang et al.
+(2021) and MPNet Song et al.
+(2020).
+This makes it a fertile ground for testing AL’s utility.
+In all this, our motivation is to not disapprove of AL as an area for research, but to motivate the inclusion of multiple practical challenges in future studies.
+Contributions: Our primary contribution is a rigorous empirical analysis of the learning behavior of AL techniques over multiple text classification pipelines, that is targeted towards answering the questions asked above.
+Additionally, we open source an AL evaluation framework [^3] Our code is available here: https://anonymous.4open.science/r/On_the_Fragility_of_Active_Learners/README.md., to enable researchers to not only reproduce our analysis, but also to rigorously evaluate their own contributions.
+2 Previous Work
+Critique of AL is not new.
+Attenberg and Provost (2011) criticize AL for its unpredictable (for a task) warm-up times, i.e., a minimum number of labeled instances before which gains over random sampling are evident.
+Margatina and Aletras (2023) point out problems with AL simulations.
+Lüth et al.
+(2023) identify key issues leading to a lack of realistic AL evaluations and propose solutions that they apply to image classification.
+Lowell et al.
+(2019) study AL empirically but focus on the interesting notion of successor models, i.e., future models that would use the labeled data collected via AL using a specific model.
+Zhan et al.
+(2021) examine the empirical effectiveness of AL, but they don’t evaluate on NLP tasks.
+Siddhant and Lipton (2018) is an empirical study of AL effectiveness similar in spirit to ours, but they focus on deep Bayesian methods.
+This work differs from from existing literature wrt being a combination of: focusing on text classification, being empirical, employing a breadth of models (traditional and deep learning based) and employing recent techniques, e.g., MPNet Song et al.
+(2020), REAL Chen et al.
+(2023).
+While some conclusions we draw here might be similar to those reported earlier, we note that it is important to revise our collective mental models in a fast evolving area such as NLP, and in enabling that, even such conclusions are valuable.
+3 Batch Active Learning - Overview
+In this work, we specifically study the batch AL setting for text classification.
+Here, a QS identifies a batch of $b$ unlabeled points, at each iteration $t$, for $T$ iterations.
+A model $M_{t}$, that is trained on the accumulated labeled pool, is produced at the end of each iteration.
+The first iteration uses a seed set of $s$ randomly sampled points (although other strategies may be used).
+We note that that $M_{t}$ should be produced using a model selection strategy (we use a hold-out set here), and must also be calibrated (we use Platt scaling Platt (2000); Niculescu-Mizil and Caruana (2005)).
+The former ensures that $M_{t}$ doesn’t overfit to the labeled data, which is likely in the initial iterations due to small quantities.
+The latter is required since many query strategies rely on uncertainty/confidence scores produced by $M_{t}$.
+Unfortunately, in our experience, multiple implementations/studies miss one or both of these steps.
+To avoid any ambiguity, we provide pseudo-code for this AL setting in Algorithm 1 in §A.
+4 Experiment Setup
+In this section, we describe our experiment setup in detail.
+4.1 Configuration Space of Experiments
+Figure 1: The space of experiments is shown.
+See §4.1 for description.
+All representations are produced by pre-trained models, which are ubiquitous in practice today.
+The lines between the boxes “Representation” and “Classifier” denote combinations that constitute our prediction pipelines.
+Note that RoBERTa is an end-to-end predictor, where there are no separate representation and classification steps.
+Our experiment configurations vary wrt datasets, text representations, classifiers, the batch and seed sizes, and of course, the QS.
+We study the following QS here: (1) Random as baseline, (2) Margin [^4] Also referred to as Smallest Margin or Breaking Ties, it is still considered to be competitive Schröder et al.
+(2022).
+Scheffer et al.
+(2001); Schröder et al.
+(2022), (3) Contrastive Active Learning (CAL) Margatina et al.
+(2021), (4) Discriminative Active Learning (DAL) Gissin and Shalev-Shwartz (2019); Ein-Dor et al.
+(2020), and (5) Representative Errors for Active Learning (REAL) Chen et al.
+(2023).
+We picked these either because they are contemporary, e.g., REAL, DAL, CAL, or have produced strong contemporary results, e.g., Margin.
+Figure 1 enumerates the configuration space.
+For further details (including hyperparameters) see §B and §E.
+Note that all representations used are based on pre-trained models which have grown quite popular in the past few years.
+For classification, we picked one each of a linear, non-linear and Deep Learning based classifier.
+Since batch or seed sizes are inconsistent in AL literature, e.g., DAL, REAL and CAL respectively use batch sizes of $50$, $150$, $2280$ we vary these settings as well.
+For an idea of the breadth of this search space, see Figure 2 which shows results for the dataset agnews and batch/seed size of $(200,200)$.
+4.2 Metrics and Other Settings
+The classifier accuracy metric we use is the F1 (macro) score, since it prevents performance wrt dominant classes from overwhelming results.
+For measuring the effectiveness of a QS,we use the relative improvement wrt the random QS of the classifier score (see Equation 1).
+The size of the unlabeled pool is $\boldsymbol{20000}$ at the start of each experiment.
+If the original dataset has more than than $20000$ instances, we extract a label-stratified sample, to retain the original class distribution.
+The size of the test set is $\boldsymbol{5000}$ also a label-stratified sample from the corresponding test set of the original dataset.
+We run an experiment till the size of the labeled set has grown to $5000$ instances [^5] Beyond this labeled set size (unrelated to the test set size) different QSes produce similar gains - see §C..
+This implies $\boldsymbol{T}=(5000-200)/200\boldsymbol{=24}$ iterations for the batch/seed size setting of $(200,200)$, and similarly $\boldsymbol{T=9}$ iterations for the $(500,500)$ setting.
+As shown in Figure 1 we have $\boldsymbol{350}$ unique configurations.
+We also execute each configuration three times in the interest of robust reporting.
+This gives us a a total of $350\times 3=\boldsymbol{1050}$ trials.
+For each AL iteration of each of these trials, we follow the due process of model selection and calibration [^6] RoBERTa is the only exception since it is naturally well-calibrated Desai and Durrett (2020)..
+Figure 2: F1 macro scores on the test set at each iteration, for the dataset agnews and batch size of $200$.
+The $x$-axes show size of the labeled data, the $y$-axes show the F1-macro scores on the test data.
+4.3 Notation and Terminology
+We introduce some notation here that will help us precisely describe our analysis in later sections.
+Let $f$ be a function that computes the model metric of interest, e.g., F1-macro. This accepts, as parameters, the random variables $h,q,d,b,s,n$, which are defined as follows:
+  • $h\in H$, the set of prediction pipelines.
+  • $q\in Q$, the set of query strategies. For convenience, we also define $q_{R}$ to be the random QS, and $\mathcal{Q}_{NR}=\{cal,dal,real,margin\}$, i.e., the subset of non-random QS.
+  • $d\in D$, the set of datasets.
+  • $(b,s)\in V$, the set of batch and seed size combinations, i.e., $V=\{(200,200),(500,500)\}$
+  • $n$ is the size of the labeled data. In our experiments, $s\leq n\leq 5000$.
+A specific value is indicated with a prime symbol on the corresponding variable, e.g., $h^{\prime}$ is a specific prediction pipeline.
+QS Effectiveness: We evaluate a non-random QS by measuring the relative improvement wrt the random QS, at a given number of labeled instances $n^{\prime}$. We use the shorthand $\delta$:
+$\delta(f(h,q,d,b,s,n^{\prime}))=100\times$
+$\frac{f(h,\boldsymbol{q},d,b,s,n^{\prime})-f(h,\boldsymbol{q_{R}},d,b,s,n^{\prime})}{f(h,\boldsymbol{q_{R}},d,b,s,n^{\prime})}$ (1)
+4.4 Decision Model
+Before looking at the results, we formalize the decision model of a practitioner using our notation. This helps us justify the aggregations we perform over results of individual experiments.
+Because of lacking prerequisite checks, there is no preference for picking a factor in combination with others.
+We model them as independent variables, i.e., the probability of a configuration is $p(h)p(q)p(d)p(b,s)$.
+Since each of these probabilities is also uniform, e.g., the general practitioner is equally likely to encounter any dataset $d\in D$, each configuration has an identical probability of occurrence [^7] They may inherit an environment with a specific prediction pipeline or a query strategy - we also present these conditional results.
+But within these conditions, the other factors are assumed to be independent and individually uniform.: $1/(|H|\times|Q|\times|D|\times|V|)$.
+In other words, any expectation we wish to compute over these settings under this decision model is a simple average.
+5 Results
+We are now ready to look at the results of our experiments.
+5.1 Expected Gains from AL
+Figure 3: Expected relative improvement in F1-macro score over random.
+(a)-(e) show this for different predictors and QS, at different training sizes (see titles).
+These correspond to Equation 2.
+(f) and (g) show marginalized improvements for different predictors and QSes respectively; see equations 3 and 4.
+Figure 3 shows the expected relative improvement, grouped in the following ways:
   1. 1.
-    In practice, where data labeling is costly, should one just spend time seeking a good AL algorithm?
-Or should they cast their net of design choices wider to take into account the classifier and representation as well?
+    Figure 3(a)-(e): These heatmaps show the expected $\delta$ at a given number of instances $n^{\prime}\in\{1000,2000,3000,4000,5000\}$.
+A cell for predictor $h^{\prime}$ and a QS $q^{\prime}\in Q_{NR}$ in the heatmap for $n^{\prime}$ training instances shows [^8] This expectation is over batch and seed sizes at given values of $n^{\prime}$; but note, different batch sizes don’t produce same values for $n^{\prime}$.
+This is explicitly reconciled - see §F.:
+    $\mathbb{E}_{d,b,s}[\delta(f(h^{\prime},q^{\prime},d,b,s,n^{\prime}))]$ (2)
+    The rows are arranged roughly in increasing order of classifier capacity, i.e., LinSVC, RF, RoBERTa, and within a group, in increasing order of approximate representation quality: word vectors (WV), USE, MPNet [^9] The relative ordering of USE vs MPNet was obtained from the Massive Text Embedding Benchmark (MTEB) rankings, where MPNET leads USE by $\sim 100$ positions today..
   2. 2.
-    An important part of assessing a technique is how we measure, i.e., metrics used.
-We discuss this aspect too, e.g., rank-based metrics such as the Wilcoxon test or Mean Rank ignore how large is the accuracy gap wrt random sampling.
-    The practical implication of this is that a small gap might not justify the cost of an AL workflow, given that random sampling is trivial to use in most programming languages.
+    Figure 3(f): This shows $\delta$ only for prediction pipelines, marginalizing over QSes.
+This is easy to show in a standard line-plot.
+The $y$-value for $x=n^{\prime}$ for predictor $h^{\prime}$ denotes:
+    $\mathbb{E}_{d,b,s,q\in Q_{NR}}[\delta(f(h^{\prime},q,d,b,s,n^{\prime}))]$ (3)
   3. 3.
-    Finally, this cautions us as consumers of research that an AL technique is not just the QS; reported results must be carefully understood to be situated within, and affected by, other significant experiment variables such as the classifier, representation, batch size, etc.
-To our knowledge, this is the first such comprehensive study of AL algorithms for the task of text classification. The rest of the paper is organized as follows:
-2 Previous Work
-The field of AL has seen a steady stream of novel contributions over the years, e.g., [16, 12, 29, 7].
-Being a hard problem - identifying important instances having only seen a few examples from the instance-label distribution - it also seems to suffer from real-world challenges, e.g., wrt variations in dataset, batch sizes, classifier used, which has been pointed out in the literature [1, 35, 18, 20].
-In the domain of NLP, the problem is compounded by the fact that text representations may vary widely.
-This is pertinent since a lot of rich representations have been proposed in the past few years, e.g., word2vec [22], Universal Sentence Encodings (USE) [3], Backpacks [14], BERT [10], MiniLM [33], MPNet [31].
-We are not aware of any work that studies AL’s effectiveness in context of these relatively modern representations.
-Hence, this work constitutes a useful addition to the line of work that draws attention to challenges of AL deployment.
-3 Batch Active Learning - Overview
-In this work, we study AL techniques for the task of text classification, in the batch AL setting. To ensure no ambiguity, we provide pseudo-code for this AL setting in Algorithm 1.
+    Figure 3(g): This is analogous to (f) and shows $\delta$ for QSes while marginalizing over predictors. The $y$-value for a specific $x=n^{\prime}$ for QS $q^{\prime}\in Q_{NR}$ denotes:
+    $\mathbb{E}_{d,b,s,h}[\delta(f(h,q^{\prime},d,b,s,n^{\prime}))]$ (4)
+Observations: In Figure 3(a)-(e), we see that as we move towards the right, the number of cells with $\delta\gtrapprox 0$ increases.
+This suggests that, in general, as the pool of labeled instances grows, AL becomes more effective.
+This might seem promising at first, but note that (a) we cannot predict when this happens in practice: we lack the theoretical tools, and it varies wrt both the predictor and the QS, and (b) if you look closely, its not that AL is becoming more effective but, rather, all configurations are converging towards [^10] This is something we observe in a separate analysis as well - see §C.
+In fact, this is the reason why we grow the labeled set to only $5000$ instances in our experiments - mentioned in §C.
+$\delta=0$.
+In other words, in low label regimes, where we expect AL to benefit us, there can be a lot of variance - it might even under-perform random sampling - and at high label regimes, their performance, even if positive, is not very different from random sampling.
+Among predictors (Figure 3(f), but this is also apparent in (a)-(e)), for RoBERTa we consistently observe $\delta>0$.
+But we note that this value isn’t high, i.e., $\delta\approx 1$.
+Among QSes, REAL and Margin, seem to do well at larger data regimes - as visible in Figure 3(g), but also in (d) and (e).
+The performance of Margin might seem somewhat surprising, since this is an old technique (proposed in Scheffer et al.
+(2001)), but similar observations have been reported elsewhere Schröder et al.
+(2022).
+5.2 Always ON Mode
+Another question we might ask is that even if AL doesn’t always surpass random, is there a downside to making it a permanent part of a labeling workflow - multiple tools allow this today [^11] Important: We have not evaluated these tools.
+They are cited as examples of common tools used in data labeling workflows in the industry., e.g., Montani and Honnibal ; Tkachenko et al.
+(2020-2022)?
+Table 1 shows some relevant numbers.
+Avg. for % times $\delta<0$ $\overline{\delta}_{\geq 0}$ $\overline{\delta}$
+Overall 51.82 0.89 -0.74
+LinSVC-WV 61.71 0.70 -1.90
+LinSVC-USE 61.57 0.46 -0.64
+LinSVC-MP 63.71 0.40 -1.48
+RF-WV 47.29 1.31 -0.30
+RF-USE 60.57 0.71 -0.63
+RF-MP 60.14 0.60 -1.24
+RoBERTa 7.71 1.29 1.01
+CAL 55.60 0.81 -1.07
+DAL 70.12 0.82 -1.29
+Margin 38.45 0.97 -0.25
+REAL 43.10 0.89 -0.34
+[…]
+Observations: In general, (first row, “Overall”), the number of incidents where the relative improvement was strictly negative (counted at various labeled data sizes across configurations) is $51.82\%$.
+This might be suggested by the heatmaps in Figure 3(a)-(e) as well, where approximately the left upper triangle of the plots combined indicates $\delta<0$.
+The average improvement when AL is as good as random is low, i.e., $\overline{\delta}_{\geq 0}=0.89$, and on the whole this quantity is actually negative, i.e., $\overline{\delta}=-0.74$.
+Again, the use of RoBERTa leads to favorable scores.
+Among QSes, Margin and REAL perform relatively well.
+Under our decision model - §4.4 - the practical implication is bleak: in the “Always ON” mode, stopping labeling early risks negative improvement.
+The only way to ensure $\delta\geq 0$ is to accumulate quite a few labels, i.e., move out of the left upper triangular region in Figure 3(a)-(e), but then the average improvement is low.
+Essentially, the “Always ON” mode is viable if the small relative gains from labeling $4000{-}5000$ instances are useful.
+5.3 Effect of Prediction Pipeline vs QS
+Papers on AL typically contribute QSes. Here we ask if that focus is warranted, i.e., what has a greater impact? - the QS or the prediction pipeline?
+We might suspect that it is the pipeline, given the performance of RoBERTa in both Figure 3 and Table 1.
+To precisely assess their relative effect, we calculate the difference in outcomes produced by changing the QS vs the pipeline.
+Here’s how we obtain such outcome data:
+  1. 1.
+    Take the example of QSes.
+For each non-random QS $q^{\prime}$, we list the scores $\delta(f(h,q^{\prime},d,b,s,n))$ for different values of $h,d,b,s,n$.
+Since there are four non-random QSes, this gives us four sets of matched observations.
+  2. 2.
+    We follow an analogous procedure for prediction pipelines, where we obtain seven matched observation sets.
+A standard method for such analysis is the Friedman test Friedman (1937), but note here that the number of matched observations for the two cases might be different, which implies different statistical power.
+Also we might not directly compare the p-values since they are a measure of significance.
+Instead we use Kendal’s W to directly measure the effect size Tomczak and Tomczak (2014).
+These effect sizes for the QS and pipeline parameters respectively are $0.34$ and $0.25$; the effect size here measures agreement, i.e., using different QSes produce similar results (higher agreement), relative to using different pipelines.
+We also built an Explainable Boosting Machine (EBM) Lou et al.
+(2013, 2012) on our observations, which is a form of Generalized Additive Model that takes into account pairwise interactions.
+The global feature importances for QS+pipeline, QS and pipeline respectively are 0.35, 0.39, 0.61 - which (a) justifies looking at the marginal effects since the importance score for QS and pipeline independently are greater than QS+pipeline, and (b) corroborates that changing pipelines has a greater impact.
+5.4 Effect of Batch/Seed Size
+We perform a Wilcoxon signed-rank test Wilcoxon (1945) to assess the effect of batch/seed sizes on $\delta$.
+This is a paired test and ideally we should match observations $\delta(f(h,q,d,200,200,n))$ and $\delta(f(h,q,d,500,500,n))$.
+However, recall that since different batch/seed sizes don’t lead to the same values of $n$ we explicitly align the sizes for such comparison (detailed in §F).
+The overall p-value of $0.90$ indicates that our batch/seed settings don’t influence $\delta$ in general.
+The exception is RoBERTa, with p-value$=1.32e{-10}$.
+A further one-sided test tells us that the batch/seed size setting of $(200,200)$ leads to greater $\delta$ values (p-value$=6.57e{-11}$).
+Predictor p-value QS p-value
+LinSVC-WV $0.18$ CAL $0.77$
+LinSVC-USE $0.41$ DAL $0.02$
+LinSVC-MP $0.60$ Margin $0.32$
+RF-WV $0.13$ REAL $0.07$
+RF-USE $0.03$
+RF-MP $0.03$
+RoBERTa $1.32e{-10}$
+Overall: $0.90$
+Table 2: The p-values for a two-sided Wilcoxon signed-rank test over $\delta$ values, from using batch/seed size $(200,200)$ vs $(500,500)$. See §5.4 for details.
+5.5 Effect of Representation
+Finally, we assess the effect of text representation on relative improvements.
+Since we want to evaluate representations alone (the prediction pipeline as a whole was already evaluated in §5.1), we ignore RoBERTa for this exercise, since its an end-to-end classifier.
+Figure 4 shows how the relative improvement $\delta$ varies with the embedding used, marginalized over other configuration variables.
+Figure 4: Effect of text representations on the relative improvement.
+We note that USE outperforms MPNet.
+This is surprising to us because on the MTEB Muennighoff et al.
+(2022) benchmarks MPNet scores much higher.
+A hypothesis that might explain both results is that USE doesn’t capture fine-grained contexts as much as MPNet does; while this might be problematic for MTEB (esp.
+tasks that rely on precise similarity measurement, such as retrieval), the fuzzier embedding space of USE is better in terms of covering the concept space in the dataset earlier in the AL process.
+This enables better assessment of informativeness, and therefore, sampling, by a non-random QS.
+6 Summary and Conclusion
+After extensive evaluation of different AL algorithms, we are forced to conclude that it is difficult to practically benefit from AL.
+Gains from QSes are inconsistent across datasets, prediction pipelines and text representations.
+In fact, between QSes and prediction pipelines, the latter seems to have a greater influence on the relative improvement over random (§5.3).
+The only general pattern we see is that positive relative improvements become likely as labeled instances accumulate; but these improvements are too small to be broadly useful (§5.1).
+Another reason as to why it is hard to derive any practical advice is that we lack the tools, theoretical or empirical, to identify a settings-specific warm-start size; when do we stop labeling to realize gains, however small?
+Further, we noted in §5.2 that using AL in an “Always ON” mode can actually perform worse than random sampling.
+The use of RoBERTa as the prediction pipeline is the only (isolated) case where we see consistent positive relative improvements.
+Our hypothesis as to why is that an end-to-end classifier has a more coherent view of the overall distribution, and therefore informativeness of a sample.
+But, obviously, we can’t discount the role that RoBERTa’s specific pre-training might play here, and further experimentation is required to disentangle their respective influences.
+Even in this case, we point out that (1) it provides further evidence for the argument that the QS is not the deciding factor, and (2) while positive, the gains aren’t considerable, with $\delta\approx 1$ (see Figure 3 (a)-(f)).
+Although extensive, this study may be considered “limited” relative to real-world variances, e.g., many more choices of classifiers, datasets, which leads us to suspect that the true picture is probably more dismal.
+What might we do to make the field of AL more useful?
+We feel the biggest problem in AL use is that practitioners have to blindly guess what specific AL technique will work best for their problem.
+As a field we need to embrace a broader discourse where the success of a technique needs to be tied to fundamental properties of datasets, e.g., topological features Chazal and Michel (2021), and predictors, e.g., VC dimension Vapnik (1995), that are identifiable in an unsupervised manner in novel settings.
+7 Limitations
+Being an empirical work, our conclusions are tied to the algorithms and settings analyzed.
+8 Acknowledgements
+We thank Sashank Gummuluri for early results that indicated that the broader area of Active Learning needs a closer investigation under practical settings.
+Joshua Selinger reviewed multiple drafts of the paper, and proposed alternative methods of measurement, for which we are grateful to him.
+We also owe a debt of gratitude to Mandar Mutalikdesai for his encouragement and continued support of this project.
+Appendix A Pseudo-code for Batch Active Learning
 Input: Unlabeled data $X_{U}$, test data $(X_{test},Y_{test})$, query strategy $\mathcal{Q}$, seed set selection strategy $\mathcal{A}$, search space $\Theta$ for model $\mathcal{M}$, seed size $s$, batch size $b$, number of iterations $T$, metric $\mathcal{V}$
@@
 7       $(X_{L,t}^{new},Y_{L,t}^{new})\leftarrow\text{obtain labels for }X_{L,t}^{new}$
-8       $(X_{L,t},Y_{L,t})\leftarrow\text{add }(X_{L,t}^{new},Y_{L,t}^{new})\text{ to previous labeled pool }(X_{L,t-1},y_{L,t-1})$
-       $M_{t}\leftarrow\operatorname*{arg\,max}_{\theta\in\Theta}M_{\theta}((X_{L,t},Y_{L,t}))$ // both model selection and calibration are performed
-9       $\mathcal{V}_{t}\leftarrow\mathcal{V}(M_{t}(X_{test}),Y_{test})$
+8       $(X_{L,t},Y_{L,t})\leftarrow\text{add }(X_{L,t}^{new},Y_{L,t}^{new})\text{ to }(X_{L,t-1},y_{L,t-1})$
+9       $M_{t}\leftarrow\operatorname*{arg\,max}_{\theta\in\Theta}M_{\theta}((X_{L,t},Y_{L,t}))$ $\mathcal{V}_{t}\leftarrow\mathcal{V}(M_{t}(X_{test}),Y_{test})$
 10       $result\leftarrow result\cup\{(\mathcal{V}_{t},t)\}$
@@
   2. 2.
-    The model is also calibrated [^1] A notable exception is in our use of the RoBERTa model, which already is well calibrated [9]..
+    The model is also calibrated [^12] A notable exception is in our use of the RoBERTa model, which already is well calibrated Desai and Durrett (2020)..
 This is critical since query strategies $\mathcal{Q}$ often use the predicted class probabilities from $M_{t}$.
-We use Platt scaling [26, 24].
+We use Platt scaling Platt (2000); Niculescu-Mizil and Caruana (2005).
 The process is initialized by selecting a seed set of size $s$ from the unlabeled data pool, using a strategy $\mathcal{A}$ (line 2). We use random selection for this step.
-We also note that a “model” here might mean a combination of a text representation, e.g., word vectors, and a classifier, e.g., Random Forest; further detailed in Section 4.
-It might seem unusual that we explicitly mention model selection and calibration, which should be part of the standard practice for building probabilistic classifiers.
-Unfortunately, in our experience multiple implementations/studies miss one or both of these steps.
-4 Comparison Methodology
+We also note that a “model” here might mean a combination of a text representation, e.g., word vectors, and a classifier, e.g., Random Forest; further detailed in Section 4.1.
+Appendix B Experiment Configurations
 In our experiments, we vary classifiers, text representations (we often jointly refer to them as a prediction pipeline), batch size, seed size and, of course, query strategies.
+These combinations are visualized in Figure 1, and are detailed in Section.
 These combinations are listed below:
@@
     1. (a)
-      Separate representation and classifier: The representations used are USE [3], MPNet [31] and word vectors (we use the models provided by the spaCy library [15]).
-While USE and MPNet directly convert sentences into vectors, in the case of word vectors, we compute the average of the vectors of all words in a text to obtain its representation.
-For classification, we useRandom Forests (RF) [2] and Support Vector Machines [6] with a linear kernel - we’ll term the latter as “LinearSVC”.
+      Separate representation and classifier: The representations used are USE Cer et al.
+(2018), MPNet Song et al.
+(2020) and word vectors [^13] The vectors of all words in a sentence are averaged to obtain its representation.
+(we use the models provided by the spaCy library Honnibal et al.
+(2020)).
+For classification, we use Random Forests (RF) Breiman (2001) and Support Vector Machines Cortes and Vapnik (1995) with a linear kernel - we’ll term the latter as “LinearSVC”.
       We use off-the-shelf representations and they are not fine-tuned on our data. Only the classifiers are trained on our data.
     2. (b)
-      End-to-end classifier: This does not require a separate representation model. We use RoBERTa [17] (a variant of BERT). This is fine-tuned on the labeled data at each AL iteration.
-    Hyperparameter search spaces are detailed in Section 0.A.4.2 of the Appendix.
+      End-to-end classifier: This does not require a separate representation model. We use RoBERTa Liu et al. (2019) (a variant of BERT). This is fine-tuned on the labeled data at each AL iteration.
+    Hyperparameter search spaces are detailed in Section E.2 of the Appendix.
 As noted in Section 3, model selection and calibration are performed during training of a prediction pipeline.
-The only exception is RoBERTa, which has been shown to be well-calibrated out of the box [9].
+The only exception is RoBERTa, which has been shown to be well-calibrated out of the box Desai and Durrett (2020).
     The first category gives us $2\times 3=6$ combinations. Counting RoBERTa, we have $\boldsymbol{7}$ prediction pipelines in our study.
@@
       Margin [^14] Also referred to as Smallest Margin or Breaking Ties.
-[27] (2001): this selects instances with the smallest differences between the confidence of the most likely and the second-most likely predicted (by the current classifier [^3] Note that in reference to Algorithm 1, at iteration $t$, the query strategy $\mathcal{Q}$ uses model $M_{t-1}$.) classes.
-Despite being a relatively old technique, it continues to be competitive [28].
+Scheffer et al.
+(2001) (2001): this selects instances with the smallest differences between the confidence of the most likely and the second-most likely predicted (by the current classifier [^15] Note that in reference to Algorithm 1, at iteration $t$, the query strategy $\mathcal{Q}$ uses model $M_{t-1}$.) classes.
+Despite being a relatively old technique, it continues to be competitive Schröder et al.
+(2022).
     3. (c)
-      Contrastive Active Learning (CAL) [21] (2021): chooses instances whose predicted class-probability distribution is the most different (based on KL divergence) from those of their $k$-nearest neighbors.
-This is similar to another work [23], where such conflicts are detected using the explanation space produced by XAI techniques.
+      Contrastive Active Learning (CAL) Margatina et al.
+(2021) (2021): chooses instances whose predicted class-probability distribution is the most different (based on KL divergence) from those of their $k$-nearest neighbors.
+This is similar to another work Nguyen and Ghose (2023), where such conflicts are detected using the explanation space produced by XAI techniques.
     4. (d)
-      Discriminative Active Learning (DAL) [13, 11] (2019): a binary classifier (a feedforward neural network) is constructed to discriminate between labeled and unlabeled data, and then selects unlabeled instances with the greatest predicted probability of being unlabeled.
+      Discriminative Active Learning (DAL) Gissin and Shalev-Shwartz (2019); Ein-Dor et al.
+(2020) (2019): a binary classifier (a feedforward neural network) is constructed to discriminate between labeled and unlabeled data, and then selects unlabeled instances with the greatest predicted probability of being unlabeled.
 This picks examples that are most different from the labeled instances in this classifier’s representation space.
-While the original work [13] only considers image datasets, a separate study shows its efficacy on text [11].
+While the original work Gissin and Shalev-Shwartz (2019) only considers image datasets, a separate study shows its efficacy on text Ein-Dor et al.
+(2020).
     5. (e)
-      Representative Errors for Active Learning (REAL) [5] (2023): identifies clusters in the unlabeled pool and assigns the majority predicted label as a “pseudo-label” to all points in it.
+      Representative Errors for Active Learning (REAL) Chen et al.
+(2023) (2023): identifies clusters in the unlabeled pool and assigns the majority predicted label as a “pseudo-label” to all points in it.
 Instances are then sampled whose predictions differ from the pseudo-label.
@@
     Datasets: we use $\boldsymbol{5}$ standard datasets: agnews, sst-2, imdb, pubmed and dbpedia-5 (a $5$-label version of the standard dbpedia dataset that we created).
-These are detailed in Table 2 in the Appendix.
+These are detailed in Table 3.
+The extent of class imbalance is represented by the label entropy column, which is calculated as $\sum_{i\in C}-p_{i}\log_{|C|}p_{i}$, with $C$ being the set of classes.
+    Dataset # classes Label entropy Description
+    sst-2 2 1.0 Single sentences extracted from movie reviews with their sentiment label Socher et al. (2013).
+    imdb 2 1.0 Movie reviews with corresponding sentiment label Maas et al. (2011).
+    agnews 4 1.0 News articles with their topic category Zhang et al. (2015).
+    pubmed 4 0.9 Sentences in medical articles’ abstracts which are labeled with their role on the abstract Dernoncourt and Lee (2017).
+    dbpedia-5 5 1.0 A subset of dbpedia Zhang et al.
+(2015) which contains Wikipedia articles accompanied by a topic label.
+The original dataset’s instances are evenly distributed across $14$ classes.
+To form dbpedia-5, we use only the first $5$ classes: Company, EducationalInstitution, Artist, Athlete, OfficeHolder.
+This was done to reduce the training time of one-vs-all classifiers, e.g., LinearSVC.
+    Table 3: Datasets used. Label entropy represents class imbalance - see §B for description.
   4. 4.
@@
     Trials: For statistical significance, we run $\boldsymbol{3}$ trials for each combination of the above settings.
-Other settings: The accuracy metric we use is the F1 (macro) score, since it prevents performance on a dominant class from overwhelming results.
-The size of the unlabeled pool is $20000$ at the start of each experiment.
-If the original dataset has more than than $20000$ instances, we extract a label-stratified sample, to retain the original class distribution.
-The size of the test set is $5000$ also a label-stratified sample from the corresponding test set of the original dataset.
-We run an experiment till the size of the labeled set has grown to $5000$ instances.
-This implies $T=(5000-200)/200=24$ iterations for the batch/seed size setting of $(200,200)$, and similarly $9$ iterations for the $(500,500)$ setting.
-Total # experiments [^4] Our code is available here: https://anonymous.4open.science/r/On_the_Fragility_of_Active_Learners/README.md.: Accounting for the above settings, we have $7\times 5\times 5\times 2=350$ unique configurations.
-Accounting for the trials, we have a total of $350\times 3=\boldsymbol{1050}$ experiments.
-Reiterating something we had mentioned earlier: we follow the due process of performing model selection and calibration (excepting RoBERTa) in each AL iteration of each of these experiments.
-5 Reproducibility of Benchmarked Techniques
-Before we analyze our results, we would like to stress that our intention is not to suggest that the techniques we evaluate, e.g., REAL, CAL, DAL, don’t work.
+Appendix C In what data regimes do query strategies most differ?
+We would intuitively expect that F1-macro scores from different QSes (for a given pipeline and dataset) should converge as we see more data due to at least two reasons:
+  • The concept space in the data would be eventually covered after a certain number of instances.
+Adding more data isn’t likely to add more information, i.e., there are diminishing returns from adding more data.
+  • At later iterations, there is less of the unlabeled pool to choose from.
+Indeed, Figure 5 confirms this.
+We first compute variances in F1-macro scores for each different pipeline/dataset combination [^16] This step comes first since the accuracies obtained by a LinearSVC would be very different from those by RoBERTa, and we don’t want to mix them.
+across QSes at a given labeled set size.
+And then we average these variances across datasets and pipelines - this is the $y$-axis.
+We see that the expected variance shrinks after a while, and at $5000$ labeled points it is close to zero, i.e., the differences from using different QSes, pipelines etc isn’t much.
+This is why we restrict the labeled set size to $5000$ instances in our experiments (as mentioned in §4.2).
+Figure 5: Expectation over variance of F1-macro given a pipeline and dataset, plotted against size of labeled data. Note that the batch/side sizes don’t strongly influence trends.
+Appendix D Reproducibility Experiments
+As mentioned earlier, our intention is not to suggest that the techniques we evaluate, e.g., REAL, CAL, DAL, don’t work.
 In the specific settings discussed in their respective papers, they most likely perform as reported.
-In the interest of fairness, we have conducted limited independent tests that confirm this - see Section 0.A.3 in the Appendix.
-6 Results
-We present our observations in this section. We begin with visualizing a cross-section of our experiments: Figure 1 shows results for the dataset agnews and batch/seed size of $(200,200)$.
-Figure 1: F1 macro scores on the test set at each iteration, for the dataset agnews and batch size of $200$.
-The $x$-axes show size of the labeled data, the $y$-axes show the F1-macro scores on the test data.
-The corresponding plot for batch size = $500$ is shown as Figure 8 in the Appendix.
-We introduce some notation to aid in the interpretation of our results. We think of the F1-macro score to be provided by this function:
-$f(g,\mathcal{Q},\mathcal{D},b,s,N)$ (1)
-Here,
-  • $g$ denotes a prediction pipeline.
-  • $\mathcal{Q}\in\{cal,dal,real,margin,random\}$ is a query strategy.
-  • $\mathcal{D}\in\{imdb,agnews,dbpedia-5,pubmed,sst-2\}$ is a dataset.
-  • $b\in\{200,500\}$ is the batch size.
-  • $s\in\{200,500\}$ is the seed size.
-  • $N$ is the size of the labeled data. Here $N\leq 5000$.
-[…]
-We now try to answer some interesting questions based on our results.
-In most cases, we will average over various factors, especially datasets, since most AL techniques assume their query strategy may be used without any prior knowledge.
-6.1 In what data regimes do query strategies most differ?
-We immediately note in Figure 1 that learning curves vary quite a bit based on the prediction pipeline.
-What broad trend(s) might we expect?
-F1-macro scores from different query strategies (for a given pipeline and dataset) should converge as we see more data due to at least two reasons: (a) the concept space would be eventually covered, (b) at later iterations, there is less of the unlabeled pool to choose from.
-Indeed, Figure 2 confirms this.
-We first compute variances in F1-macro scores for different pipeline/dataset combinations [^5] This step comes first since the accuracies obtained by a LinearSVC would be very different from those by RoBERTa.
-In this case, the inner quantity is conditioned over the dataset first, because we wanted to validate a specific property.
-In subsequent analyses, we will average over datasets, so as to mimic real-life use of AL, where no prior knowledge is assumed., and then plot their averages.
-In other words, a point $(x,y)$ for the batch/seed size $(200,200)$ curve denotes:
-$y=\mathbb{E}_{g,\mathcal{D}}[Var[f|g=g^{\prime},\mathcal{D}=d^{\prime},N=x,b=200,s=200]]$ (2)
-where $g^{\prime}$ and $d^{\prime}$ are a specific pipeline and dataset respectively.
-Figure 2: Expectation over variance of F1-macro given a pipeline and dataset, plotted size of labeled data. See Section 6.1 for details. Note that the batch size doesn’t strongly affect trends.
-6.2 Measuring effectiveness
-A key reason to use AL is its effectiveness over random sampling.
-A common way to assess this is to use the Wilcoxon signed-rank test, e.g., [11], by comparing accuracy scores from these strategies, at various labeled data sizes.
-Unfortunately, this entirely ignores the magnitude of the difference in scores.
-This is a limitation of rank based metrics in general [^6] Another example is Mean Rank over AUC values of the learning curves..
-See Figure 3, which shows learning curves for the random strategy, and query strategies $QS1$ and $QS2$.
-The latter two strategies have identical p-values of $9.77\times 10^{-4}$ for the Wilcoxon test (compared to random).
-Figure 3: Hypothetical query strategies $QS1$ and $QS2$ shown against random. Note that p-value from the Wilcoxon test is identical for both. Discussed in Section 6.2.
-In practice, one would clearly prefer $QS2$ over $QS1$, since it reaches higher accuracies earlier.
-In fact, considering the possibly non-zero cost of setting up an AL-based tagging system, a query strategy needs to be significantly better for it to be useful.
-We recommend assessing the magnitude (relative to the random strategy) using Root Mean Square Error (RMSE) or the Area Under the (Learning) Curve (AUC).
-RMSE vs AUC: the choice between them comes down to these factors: (a) RMSE doesn’t provide a direction of change (while AUC does), which may be addressed by pairing it with the Wilcoxon test, (b) AUC monotonically increases from left-to-right on the $x$-axis (unlike RMSE which is a mean), and so it must be carefully interpreted.
-Direct comparisons of AUC are most convenient over identical spans of the $x$-axis.
-We apply these metrics to our results to obtain Figure 4, where we show both average p-values from the Wilcoxon test (on the left) and average RMSE scores (right subplot).
-Mathematically, for a given value on the $x$-axis, we compare the sequences:
-$(\mathbb{E}_{g,\mathcal{D}}[f|\mathcal{Q}\in S_{qs},N=n,b=200,s=200])_{n\leq x}\text{ and}$
-$(\mathbb{E}_{g,\mathcal{D}}[f|\mathcal{Q}=random,N=n,b=200,s=200])_{n\leq x}$ (3)
-where $S_{qs}=\{cal,dal,real,margin\}$. Note that the $y$ values are cumulative, i.e., all sizes $n\leq x$ are considered. We prefer the RMSE here to normalize for the $x$-axis span.
-For the Wilcoxon test, we have set the minimal sample size at $3$ points, which is why we see different starting points on the $x$-axes in Figure 4.
-This test is also performed to assess if the alternative, i.e., a non-random strategy, produces higher values than the random strategy.
-Figure 4: The left plot shows averaged p-values from a Wilcoxon test, while the right plot shows averaged RMSE values.
-A consistently good query strategy should have a low p-value and high RMSE; none of the evaluate strategies fall in that category.
-Discussed in Section 6.2.
-An ideal query strategy should have a low p-value and high RMSE.
-Unfortunately, no query strategy fulfills this criteria: while cal and dal have high RMSE values, they have poor p-values (the dal curve eclipses cal in the plot).
-And while real and margin have p-values that improve with labeled data, they exhibit small RMSE scores.
-Some things of note:
-  1. 1.
-    The fact the p-value scores of real and margin improve with size, is something we will encounter again (Figure 5).
-  2. 2.
-    The RMSE plot reinforces trends of the variance plot in Figure 2: we see higher RMSEs in the low data regime.
-  3. 3.
-    The batch/seed settings preserve the general trend.
-6.3 Gain over random and margin strategies
-We now ask if we actually do better than random sampling. We define the relative improvement $\delta$ between two scores as:
-$\delta(treatment,control)=100\times\frac{(treatment-control)}{control}$ (4)
-For a labeled dataset of size $(n_{1},n_{2}]$, batch/seed size $(b^{\prime},s^{\prime})$ and
-$q^{\prime}\in\{cal,dal,real,margin\}$, we compute:
-$\mathbb{E}_{g,\mathcal{D}}[\delta((f|\mathcal{Q}=q^{\prime},b=b^{\prime},s=s^{\prime},n_{1}<N\leq n_{2}]),\\ (f|\mathcal{Q}=random,b=b^{\prime},s=s^{\prime},n_{1}<N\leq n_{2}))]$ (5)
-This is shown in Figure 5, where the labeled data size is broken up into multiple bins to distinguish patterns at different size regimes.
-We also separate the batch sizes (see legend).
-Given the negative scores, it seems that in general, random is either competitive or better!
-At relatively higher data sizes, margin and real start to surpass random: which seems to suggest that (a) at low size regimes, where we expect the greatest impact from AL, their performance is actually poor, and (b) after gathering some data, these strategies are able to better exploit structure in the data (and margin seems surprisingly competitive in this regime).
-Also, recall that we had seen this behaviour of margin and real in the RMSE subplot of Figure 4.
-Since behaviour across the two batch/seed size settings don’t seem to appreciably differ (we note this in Figures 2, 4 and 5), we will average over them in the next few plots.
-Figure 5: Average relative improvement in F1-macro score over random, for different query strategies. Corresponds to Equation 5.
-We can repeat the above exercise, to assess the effect of pipelines now, i.e., for a given data size bin $(n_{1},n_{2}]$ and prediction pipeline $g^{\prime}$, we now compute:
-$\mathbb{E}_{\mathcal{Q}\in S_{qs},\mathcal{D},b,s}[\delta((f|g=g^{\prime},n_{1}<N\leq n_{2}]),\\ (f|g=g^{\prime},\mathcal{Q}=random,n_{1}<N\leq n_{2}))]$ (6)
-where $S_{qs}=\{cal,dal,real,margin\}$. This is plotted in Figure 6. Here we note that, using RoBERTa gives us a better average performance over random at any size regime.
-Figure 6: Average relative improvement in F1-macro score over random, for different pipelines. This corresponds to Equation 6.
-We might want to know if for the promising case of RoBERTa relatively new techniques like dal, real and cal outperform margin. This is easy to plot with minor changes to the quantity in Equation 6:
-$\mathbb{E}_{\mathcal{Q}\in S_{qs},\mathcal{D},b,s}[\delta((f|g=RoBERTa,n_{1}<N\leq n_{2}]),\\ (f|g=RoBERTa,\mathcal{Q}=margin,n_{1}<N\leq n_{2}))]$ (7)
-where $S_{qs}=\{cal,dal,real\}$. The corresponding plot is shown in Figure 7. We see that aside from one data regime, $(1250-2500]$ margin remains quite competitive.
-Figure 7: Average relative improvement in F1-macro score over margin for RoBERTa. This corresponds to Equation 7.
-6.4 The best combinations
-Finally, we ask what are the best combinations of pipelines and query strategies, at different size regimes?
-We measure this by the AUC per regime.
-Given a pipeline $g^{\prime}$, strategy $q^{\prime}$, and data sizes $n_{i}$ within a bin, the AUC [^7] Using AUC is reasonable since the $x$-axis is broken up into bins of roughly equal size (see discussion in Section 6.2), and we are not looking for comparison to the random strategy.
-is computed for the following sequence:
-$(\mathbb{E}_{\mathcal{D},b,s}[f|g=g^{\prime},\mathcal{Q}=q^{\prime},N=n_{i}])_{n_{i}}$ (8)
-Table 1 shows the top-$10$ combinations of pipelines and query strategies for different regimes.
-Not surprisingly, wrt pipelines, RoBERTa shows the strongest performance.
-Somewhat surprisingly, the next best pipeline is LinearSVC-MPNet (shortened to LinSVC_MPNet) and not a RF-based pipeline.
-Our guess as to why this is the case is that RF’s hyperparam.
-search space (detailed in Section 0.A.4.2) was not comprehensive enough, while MPNet transformed the data to be nearly linearly separable, which enabled the high-bias LinearSVC to quickly find a good optima.
-Sadly, no single query strategy dominates the QS column, and thus we are offered no insight as to what strategy pairs best with a pipeline.
-Somewhat noteworthy is the fact that the random strategy is quite competitive for the LinearSVC-MPNet pipeline, but not for RoBERTa.
-Table 1: The top-$10$ combinations of pipelines and query strategies are shown (in two columns) for different labeled data size regimes.
-We show the AUC score for a combinations in size regime.
-See Section 6.4 for details.
-train size pipeline QS AUC pipeline QS AUC
-(0-1250] 1. RoBERTa dal 875.46 6. LinSVC_MPNet random 807.04
-2. RoBERTa real 873.64 7. LinSVC_MPNet real 794.69
-3. RoBERTa margin 873.36 8. LinSVC_MPNet margin 793.08
-4. RoBERTa cal 871.92 9. LinSVC_MPNet cal 790.42
-5. RoBERTa random 862.99 10. LinSVC_MPNet dal 789.60
-(1250-2500] 1. RoBERTa cal 986.22 6. LinSVC_MPNet random 913.76
-2. RoBERTa dal 984.40 7. LinSVC_MPNet real 907.98
-3. RoBERTa real 981.07 8. LinSVC_MPNet margin 907.39
-4. RoBERTa margin 976.07 9. LinSVC_MPNet dal 900.13
-5. RoBERTa random 974.04 10. LinSVC_MPNet cal 899.12
-(2500-3750] 1. RoBERTa margin 903.60 6. LinSVC_MPNet random 837.91
-2. RoBERTa dal 900.93 7. LinSVC_MPNet margin 836.57
-3. RoBERTa real 900.83 8. LinSVC_MPNet real 833.95
-4. RoBERTa cal 897.16 9. LinSVC_MPNet dal 828.73
-5. RoBERTa random 891.58 10. LinSVC_MPNet cal 827.24
-(3750-5000] 1. RoBERTa real 1087.92 6. LinSVC_MPNet margin 1016.94
-2. RoBERTa margin 1087.37 7. LinSVC_MPNet random 1010.88
-3. RoBERTa dal 1085.06 8. LinSVC_MPNet real 1008.67
-4. RoBERTa cal 1076.79 9. LinSVC_MPNet dal 1004.44
-5. RoBERTa random 1075.20 10. LinSVC_MPNet cal 999.97
-7 Summary and Conclusion
-We first summarize our observations:
-  1. 1.
-    We noted in the dataset cross-section (Figure 1) of our experiments that the behavior of query strategies vary widely wrt pipelines.
-However, we observe the larger expected trend of reduction in the variance of accuracies produced by different strategies, with increase in the size of labeled data (Figure 2 in Section 2).
-  2. 2.
-    Wrt assessing the value of a query strategy, we noted that the Wilcoxon test only provides a partial picture by ignoring magnitudes (Figure 3 in Section 6.2).
-We might use the RMSE score to measure differences wrt the random strategy.
-Applied to our results, it shows that no query strategy exhibits the favorable properties of having both a low p-value and high RMSE score (Figure 4 in Section 6.2).
-  3. 3.
-    Our batch/seed size settings of $(200,200)$ and $(500,500)$ only seem to influence slight change in behaviors. We noted this in Figure 2, Figure 4 and Figure 5.
-  4. 4.
-    Most query strategies exhibit no significant gain over the random strategy (Figure 5), although at relatively higher data regimes we noted real and margin surpass random.
-  5. 5.
-    Among pipelines, RoBERTa exhibits the strongest performance (Figure 6) relative to random, but not wrt margin (Figure 7).
-  6. 6.
-    When we list the top-$10$ best performing combinations per data size regime (Table 1), we note that RoBERTa and LinearSVC-MPNet consistently do well, but unfortunately, their best performances are paired with different query strategies.
-The only (broad) takeaways are that while LinearSVC-MPNet works well with the random strategy, RoBERTa performs well with non-random strategies, but its not clear which specific one.
-After extensive evaluation of different AL algorithms, we are forced to conclude that AL doesn’t reliably work across datasets, prediction pipelines and text representations.
-This study may be considered “limited” relative to real-world variances, e.g., many more choices of classifiers, datasets, which leads us to suspect that the true picture is probably more dismal.
-Of course, we see some associations in isolated pockets, such as RoBERTa consistently performs well, that are promising, but not enough (with what query strategy?) to derive a useful prescription.
-It’s known that query strategies require a “warm-start” labeled set, and this is probably why we observe the deferred performance of real and margin in Figure 5, but again, this does not translate to practical advice unless we can precisely identify a settings-specific warm-start size.
-It’d seem the only way to know if an AL technique would work in a setting is to find results reported for similar settings, and even this advice leaves open the question of inferring similarity between datasets.
-And if one can’t find such results, the only other way to know is by trying out multiple techniques - which of course is a chicken-and-egg problem when data-labeling is expensive!
-How do we encourage the field of AL to be more utility-driven?
-We believe there are at least two ways: (a) embrace a broader discourse where the success of a technique needs to be tied to fundamental properties of datasets (e.g., topological features [4]) and predictors (e.g., VC dimension [32]), that are identifiable in an unsupervised manner in novel settings, and (b) on the empirical front, reported results need to account for an array of settings much more diverse than is customary today.
-Appendix 0.A Appendix
-0.A.1 Datasets
-As an addendum to the methodology discussion in Section 4, we provide details of the datasets used in Table 2.
-Table 2: Datasets used.
-Dataset # classes Description
-sst-2 2 Single sentences extracted from movie reviews with their sentiment label [30].
-imdb 2 Movie reviews with corresponding sentiment label [19].
-agnews 4 News articles with their topic category [36].
-pubmed 4 Sentences in medical articles’ abstracts which are labeled with their role on the abstract [8].
-dbpedia-5 5 A subset of dbpedia [36] which contains Wikipedia articles accompanied by a topic label.
-The original dataset’s instances are evenly distributed across $14$ classes.
-To form dbpedia-5, we use only the first $5$ classes: Company, EducationalInstitution, Artist, Athlete, OfficeHolder.
-This was done to reduce the training time of one-vs-all classifiers, e.g., LinearSVC.
-0.A.2 Learning curves for agnews
-This is the plot for the batch/seed size setting of $(500,500)$, for the dataset agnews. This is a companion to the plot shown in Figure 1 of Section 6.
-Figure 8: F1 macro scores on the test set at each iteration, for the dataset agnews and batch size of $500$.
-The $x$-axes show size of the labeled data, the $y$-axes show the F1-macro scores on the test data.
-0.A.3 Reproducibility Experiments
+In the interest of fairness, we have conducted limited independent tests that confirm this.
 In all cases, we have attempted to replicate the original settings, e.g.
@@
 For DAL, we use the dataset cola [^17] https://nyu-mll.github.io/CoLA/ and utilise the Hugging Face library to finetune BERT (while the original work employs TensorFlow [^18] https://www.tensorflow.org/, but we use equivalent settings).
-Figure 9 shows a comparison between our results and the reported ones in these papers [21, 5, 11] for CAL, REAL, DAL, respectively.
+Figure 6 shows a comparison between our results and the reported ones in these papers Margatina et al.
+(2021); Chen et al.
+(2023); Ein-Dor et al.
+(2020) for CAL, REAL, DAL, respectively.
 Despite some minor differences in the setups, we observe that these AL methods work as described in their respective papers in these settings.
@@
 Moreover, in some cases, there is no model selection performed, which we remedy in our experiments (Section 3).
-Figure 9: Comparison between published results in [21, 5, 11] and ours with the same settings for CAL, REAL, DAL.
-Table 3: Settings for reproducibility experiments.
+Figure 6: Comparison between published results in Margatina et al. (2021); Chen et al. (2023); Ein-Dor et al. (2020) and ours with the same settings for CAL, REAL, DAL.
 AL Dataset AL loop Classifier & text representation QS parameters Metric
@@
 # clusters=25 F1-macro
-0.A.4 Hyperparameters
-0.A.4.1 Query Strategy (QS) hyperparameters
+Table 4: Settings for reproducibility experiments.
+Appendix E Hyperparameters
+E.1 Query Strategy (QS) hyperparameters
 For each QS’s hypeparameters, we use the values recommended by the authors in corresponding papers.
 This means setting number of nearest neighbors in CAL to 10, number of clusters in DAL to 25, and keeping the same discriminative model in REAL.
-0.A.4.2 Hyperparameters search for prediction pipelines
+E.2 Hyperparameters search for prediction pipelines
 Table 5 shows the search space for hyperparameters we use for each classifier.
-Table 4: Hyperparameters for each classifier in the prediction pipelines.
 Classifier Hyperparameters
@@
 max features = sqrt
+Table 5: Hyperparameters for each classifier in the prediction pipelines.
+Appendix F Averaging over Different Batch-Sizes
+When computing expectations over different batch/seed sizes (like in Equation 2) a challenge is that different settings don’t lead to same number of instances.
+For ex., for $b=200,s=200$, the size of the trained pool assumes the values 200,400,..,5000200,400,..,5000200 , 400 , .
+.
+, 5000, and for $b=500,s=500$, the sizes are 500,1000,..,5000500,1000,..,5000500 , 1000 , .
+.
+, 5000.
+To compute an expectation of the form $\mathbb{E}_{b,s}[.,n^{\prime}]$, we use the sizes from the larger batch, i.e., $n^{\prime}\in\{500,1000,..,5000\}$, and map the closest sizes from the smaller batch to them.
+For ex., here are some size mappings from the small batch case to the larger one: $800\rightarrow 1000,1000\rightarrow 1000,1200\rightarrow 1000,1400\rightarrow 1500,1600\rightarrow 1500$.
+Appendix G Always ON Mode
+Table 6 presents standard deviations for the “Always ON” case, and is a companion to Table 1 in §5.2.
+Note the extremely high variances in moving across combinations of the configurations and size of the labeled set.
+Avg. for % times $\delta<0$ $\overline{\delta}_{\geq 0}$ $\overline{\delta}$
+Overall 51.82 0.89 $\pm$ 0.92 -0.74 $\pm$ 3.02
+LinSVC-WV 61.71 0.70 $\pm$ 0.60 -1.90 $\pm$ 3.94
+LinSVC-USE 61.57 0.46 $\pm$ 0.49 -0.64 $\pm$ 1.85
+LinSVC-MP 63.71 0.40 $\pm$ 0.44 -1.48 $\pm$ 3.53
+RF-WV 47.29 1.31 $\pm$ 1.01 -0.30 $\pm$ 2.63
+RF-USE 60.57 0.71 $\pm$ 0.69 -0.63 $\pm$ 1.85
+RF-MP 60.14 0.60 $\pm$ 0.55 -1.24 $\pm$ 3.59
+RoBERTa 7.71 1.29 $\pm$ 1.17 1.01 $\pm$ 1.94
+cal 55.60 0.81 $\pm$ 0.86 -1.07 $\pm$ 3.23
+dal 70.12 0.82 $\pm$ 0.94 -1.29 $\pm$ 3.22
+margin 38.45 0.97 $\pm$ 0.88 -0.25 $\pm$ 2.78
+real 43.10 0.89 $\pm$ 0.99 -0.34 $\pm$ 2.67
+[…]

### v6

@@
 1 Introduction
-Within a supervised learning setup, Active Learning (AL) techniques use a Query Strategy (QS) to identify an unlabeled set of instances which is optimal in the following sense: if labelled and added to the training data, they lead to the greatest improvement in model accuracy, relative to any other same-sized set.
+Within a supervised learning setup, Active Learning (AL) techniques Settles (2009) use a Query Strategy (QS) to identify an unlabeled set of instances which is optimal in the following sense: if labelled and added to the training data, they lead to the greatest improvement in model accuracy, relative to any other same-sized set.
 In cases where labelling is expensive, the value proposition of AL is that it is cost-efficient compared to random sampling, and a model reaches greater accuracy with a smaller number of labelled instances.
@@
 This trust extends to related decisions such as batch and seed sizes, as well as the hyperparameters (if any) of the AL technique since there is no way to empirically pick them: to compare with random sampling, or among techniques, labels are required.
-But if one had labels, they wouldn’t need AL!
-In this sense, the AL setup is unforgiving as one needs to make the optimal choice in one shot.
-This leads us to question the validity of the implicit but consequential assumption of task transfer.
-A related question is if it makes sense to use AL in an “Always ON” mode in a data labeling workflow; this is akin to asking if AL might perform worse than random sampling.
-We need to quantify both the frequency and magnitude of gains from AL, to be able to evaluate the cost of such pipeline; even simple AL techniques require a model to be evaluated over the unlabeled data pool, which can be expensive depending on the model complexity, size of the data pool and the latency allowed per AL iteration.
+But if one had labels, they wouldn’t need AL Attenberg and Provost (2011)!
+In this sense, the AL setup is unforgiving as one needs to make the optimal choice in one shot Margatina and Aletras (2023).
+This leads us to ask multiple questions about the broader area.
+How valid is the implicit but consequential assumption of transferability?
+A related question is whether the focus on QSes alone is warranted - how much do the other components of a prediction pipeline affect outcomes?
+And finally, does it make sense to use AL at least in an “Always ON” mode in a data labeling workflow; this is akin to asking if AL might perform worse than random sampling.
+We need to quantify both the frequency and magnitude of gains from AL, to be able to evaluate the cost of such pipelines.
+This is because even simple AL techniques require a model to be evaluated over the unlabeled data pool, which can be expensive depending on the model complexity, size of the data pool and the latency allowed per AL iteration.
 To be clear, we don’t question if AL results are reproducible within the original setups they were reported in [^2] In the interest of fairness, we conducted limited reproduciblity tests for the AL techniques we benchmark here, and were able to replicate reported results - see §D.
@@
 This makes it a fertile ground for testing AL’s utility.
-In all this, our motivation is to not disapprove of AL as an area for research, but to motivate the inclusion of multiple practical challenges in future studies.
+In all this, our motivation is not to disapprove of AL as an area for research, but to motivate the inclusion of multiple practical challenges in future studies.
 Contributions: Our primary contribution is a rigorous empirical analysis of the learning behavior of AL techniques over multiple text classification pipelines, that is targeted towards answering the questions asked above.
-Additionally, we open source an AL evaluation framework [^3] Our code is available here: https://anonymous.4open.science/r/On_the_Fragility_of_Active_Learners/README.md., to enable researchers to not only reproduce our analysis, but also to rigorously evaluate their own contributions.
+Additionally, we open source an AL evaluation framework [^3] Our framework, ALchemist, is available here: https://github.com/ThuongTNguyen/ALchemist., to enable researchers to not only reproduce our analysis, but also to rigorously evaluate their own contributions.
 2 Previous Work
@@
 Siddhant and Lipton (2018) is an empirical study of AL effectiveness similar in spirit to ours, but they focus on deep Bayesian methods.
+Prabhu et al.
+(2019) study sampling biases in deep AL, but their study is limited to one prediction model - FastText.zip Joulin et al.
+(2016) - and considers only QSes based on uncertainty sampling.
 This work differs from from existing literature wrt being a combination of: focusing on text classification, being empirical, employing a breadth of models (traditional and deep learning based) and employing recent techniques, e.g., MPNet Song et al.
@@
 Note that RoBERTa is an end-to-end predictor, where there are no separate representation and classification steps.
+Also note that the popular Transformer architecture Vaswani et al.
+(2017) is represented by RoBERTa and MPNet here.
 Our experiment configurations vary wrt datasets, text representations, classifiers, the batch and seed sizes, and of course, the QS.
@@
 Note that all representations used are based on pre-trained models which have grown quite popular in the past few years.
-For classification, we picked one each of a linear, non-linear and Deep Learning based classifier.
+For classification, we picked one each of a linear, non-linear and Deep Learning based classifier [^5] Although end-to-end classifiers, e.g., RoBERTa, DistilBERT Sanh et al.
+(2020), are popular today, we include pipelines with separate representation and classification components since they are still used where: (a) a good latency-accuracy trade-off is needed, and (b) there are multiple downstream tasks that might leverage the representation, e.g., classification, similarity-based retrieval, sentiment analysis.
+On a different note, the growing popularity of Retrieval Augmented Generation (RAG) Lewis et al.
+(2020) has re-shifted focus to the area of learning good embeddings..
 Since batch or seed sizes are inconsistent in AL literature, e.g., DAL, REAL and CAL respectively use batch sizes of $50$, $150$, $2280$ we vary these settings as well.
@@
 This gives us a a total of $350\times 3=\boldsymbol{1050}$ trials.
-For each AL iteration of each of these trials, we follow the due process of model selection and calibration [^6] RoBERTa is the only exception since it is naturally well-calibrated Desai and Durrett (2020)..
+For each AL iteration of each of these trials, we follow the due process of model selection [^7] Margatina and Aletras (2023) point out that this is lacking in most AL studies.
+This is another way the current work differentiates itself.
+and calibration [^8] RoBERTa is the only exception since it is naturally well-calibrated Desai and Durrett (2020)..
 Figure 2: F1 macro scores on the test set at each iteration, for the dataset agnews and batch size of $200$.
@@
 We introduce some notation here that will help us precisely describe our analysis in later sections.
-Let $f$ be a function that computes the model metric of interest, e.g., F1-macro. This accepts, as parameters, the random variables $h,q,d,b,s,n$, which are defined as follows:
+Let $f$ be a function that computes the model metric of interest, e.g., F1-macro.
+This accepts, as parameters, the random variables [^9] Of course, in this work we consider them to only assume specifically chosen values, e.g., RF, LinSVC and RoBERTa as predictors.
+$h,q,d,b,s,n$, which are defined as follows:
   • $h\in H$, the set of prediction pipelines.
@@
 (2013, 2012) on our observations, which is a form of Generalized Additive Model that takes into account pairwise interactions.
-The global feature importances for QS+pipeline, QS and pipeline respectively are 0.35, 0.39, 0.61 - which (a) justifies looking at the marginal effects since the importance score for QS and pipeline independently are greater than QS+pipeline, and (b) corroborates that changing pipelines has a greater impact.
+The global feature importance [^15] The EBM was constructed using different train-test splits, hence both mean and standard deviations across these splits are reported.
+for QS+pipeline, QS and pipeline respectively are $0.41\pm 0.01$, $0.42\pm 0.02$, $0.63\pm 0.01$ which (a) justifies looking at the marginal effects since the importance score for QS and pipeline independently are at least as large as QS+pipeline, and (b) corroborates that changing pipelines has a greater impact.
 5.4 Effect of Batch/Seed Size
@@
 But, obviously, we can’t discount the role that RoBERTa’s specific pre-training might play here, and further experimentation is required to disentangle their respective influences.
-Even in this case, we point out that (1) it provides further evidence for the argument that the QS is not the deciding factor, and (2) while positive, the gains aren’t considerable, with $\delta\approx 1$ (see Figure 3 (a)-(f)).
+Even in this case, we point out that (1) it provides further evidence for the argument that the QS alone does not decide outcomes, and (2) while positive, the gains aren’t considerable, with $\delta\approx 1\%$ (see Figure 3 (a)-(f)).
 Although extensive, this study may be considered “limited” relative to real-world variances, e.g., many more choices of classifiers, datasets, which leads us to suspect that the true picture is probably more dismal.
@@
 Being an empirical work, our conclusions are tied to the algorithms and settings analyzed.
+In particular, the experiments (a) don’t include Large Language Models, or (b) is not exhaustive wrt hyperparameters such as batch and seed set sizes; we use two settings, but please note that there is no standard way to select values for these a priori.
+Another aspect that is not considered here is the difference between academic vs real-world datasets, which might lead to different behaviors for a QS Margatina and Aletras (2023)
+We also point out that the shortcomings of individual QSes themselves are not a limitation of this study, which maybe seen as an empirical survey with the goal of thoroughly evaluating existing QSes as-is.
 8 Acknowledgements
-We thank Sashank Gummuluri for early results that indicated that the broader area of Active Learning needs a closer investigation under practical settings.
+We thank Sashank Gummuluri for early results in various practical settings.
 Joshua Selinger reviewed multiple drafts of the paper, and proposed alternative methods of measurement, for which we are grateful to him.
 We also owe a debt of gratitude to Mandar Mutalikdesai for his encouragement and continued support of this project.
+The authors would also like to thank reviewers from the ACL Rolling Review process, especially YZn8, who helped improve various technical aspects of the paper.
 Appendix A Pseudo-code for Batch Active Learning